### v2

# Quantitative Certification of Bias in Large Language Models

Abstract
Warning: This paper contains model outputs that are offensive in nature. Large Language Models (LLMs) can produce biased responses that can cause representational harms. However, conventional studies are insufficient to thoroughly evaluate LLM bias, as they can not scale to large number of inputs and provide no guarantees. Therefore, we propose the first framework, QuaCer-B (Quantitative Certification of Bias) that certifies LLMs for bias on distributions of prompts. A certificate consists of high-confidence bounds on the probability of unbiased LLM responses for any set of prompts mentioning various demographic groups, sampled from a distribution. We illustrate the bias certification for distributions of prompts created by applying varying prefixes drawn from a prefix distributions, to a given set of prompts. We consider prefix distributions for random token sequences, mixtures of manual jailbreaks, and jailbreaks in the LLM’s embedding space to certify bias. We obtain non-trivial certified bounds on the probability of unbiased responses of SOTA LLMs, exposing their vulnerabilities over distributions of prompts generated from computationally inexpensive distributions of prefixes.
1 Introduction
Text-generating Large Language Models (LLMs) are recently being deployed in user-facing applications, such as chatbots (Lee et al., 2023). LLM-powered chatbots, like ChatGPT (Brown et al., 2020a) and Perplexity AI (Perplexity, 2023), are popular for their ability to produce human-like texts (Shahriar and Hayawi, 2023). The underlying LLMs are safety-trained (Wang et al., 2023) to avoid generating harmful content. However, despite safety training, they have been shown to produce texts that exhibit social biases and stereotypes (Kotek et al., 2023; Manvi et al., 2024; Hofmann et al., 2024). Such texts can result in representational harms (Suresh and Guttag, 2021; Blodgett et al., 2020) to protected demographic groups (a subset of the population that is negatively affected by bias). Representational harms include stereotyping, denigration, and misrepresentation of historically and structurally oppressed demographic groups. Although “representational harms are harmful in their own right” (Blodgett et al., 2020), as they can socially impact individuals and redefine social hierarchies, the resulting allocation harms can lead to economic losses to the protected groups and are hence regulated by anti-discrimination laws such as (Sherry, 1965). Language is considered an important factor for labeling, modifying, and transmitting beliefs about demographic groups and can seldom result in new social hierarchies and amplify representational harm. Hence, with the rising popularity of LLMs, it is important to formally evaluate their biases to effectively mitigate representational harms due to them.
Prior work has primarily focused on benchmarking the performance of LLMs (Liang et al., 2023; Wang et al., 2024; Mazeika et al., 2024) and adversarial attack generation (Sheng et al., 2020; Zou et al., 2023; Vega et al., 2023; Wallace et al., 2019). While these methods provide some empirical insights into LLM bias, they have several fundamental limitations (McIntosh et al., 2024; Yang et al., 2023) such as — (1) Limited test cases: Benchmarking consists of evaluating several but limited number of test cases. Due to its enumerative nature, benchmarking can not scale to the prohibitively large sets of prompts that can elicit bias from LLMs. Adversarial attacks, on the other hand, identify only a few worst-case examples of bias, which do not inform about the overall biases from large input sets; (2) Test set leakage: LLMs may have been trained on the popular benchmarking datasets, thus resulting in biased evaluation; (3) Lack of guarantees. Benchmarking involves empirical estimation that does not provide any formal guarantees of generalization over any input sets. Similarly, adversarial attacks give limited insights as they can show existence of problematic behaviors on individual inputs but do not quantify the risk of obtaining biased LLM responses.
This work. In this work, we propose an alternative to benchmarking and adversarial attacks — certifying LLMs for bias which gives formal guarantees. Certification operates on a prohibitively large set of inputs, represented as a specification. As specifications define inputs mathematically through operators over the vocabulary of LLMs, certification can provide guarantees on the behavior of the target model that generalize to unseen inputs satisfying the specification. With guarantees, we can be better informed about the available models before deploying them in public-facing applications.
Key challenges. (1) There are no existing precise mathematical representations of large sets of prompts to make practical specifications. (2) State-of-the-art neural network certifiers (Wang et al., 2021; Singh et al., 2019) currently do not scale to LLMs as they require white-box access to the model and lose precision significantly for larger models, resulting in inconclusive results.
Our approach. Given the diversity of prompts handled by LLMs, there will always be a few cases where the LLM output will be biased (e.g., these can be found with adversarial attacks (Zou et al., 2023; Chao et al., 2023)). Hence, we believe that LLM certification must be quantitative (Li et al., 2022a; Baluta et al., 2021) and focus on the following question:
What is the probability of LLM responses being unbiased for any random set of prompts each mentioning different demographic groups?
Figure 1: (Overview of QuaCer-B): QuaCer-B is a quantitative certification framework to certify the bias in the responses of a target LLM for a random set of prompts that differ by their sensitive attribute. In specific instantiations, QuaCer-B samples a (a) set of prefixes from a given distribution and prepends them to a prompt set to form (b) the prompts given to the target LLM. (c) The target LLM’s responses are checked for bias by a bias detector, (d) whose results are fed into a certifier. (e) The certifier computes bounds on the probability of obtaining biased responses from the target LLM for any set of prompts formed with a random prefix from the distribution.
We study bias over a set of prompts with varying demographic groups, inspired from Counterfactual Fairness (Kusner et al., 2018). In particular, we identify the biases across the LLM responses for counterfactual prompts (that differ only by the mentioned demographic groups) to capture the difference in their outputs caused by just varying demographic groups in the prompts. Exactly computing the probability of unbiased responses is infeasible due to the large number of possible sets of counterfactual prompts over which the biased behavior has to be determined. One can try to compute deterministic lower and upper bounds on the probability (Berrada et al., 2021). However, this is expensive and requires white-box access making it not applicable to popular, SOTA but closed-source LLMs such as GPT-4 (Achiam et al., 2023). Therefore, we focus on black-box probabilistic certification that estimates the probability of unbiased responses with high confidence bounds. We develop the first general specification and certification framework, QuaCer-B [^1] Quantitative Certification of Bias for LLM bias, applicable to both open and closed-source LLMs. Figure 1 gives an overview of our framework. Our specifications encode the probability of unbiased LLM responses for counterfactual prompts. We demonstrate QuaCer-B with $3$ exemplar kinds of specifications, each characterized by a specific kind of distribution of prefixes applied to a fixed set of counterfactual prompts. The distributions of prefixes we present are — random sequence of tokens, mixture of popular jailbreaks, and jailbreak perturbations in embedding space. The first two are model-agnostic specifications and hence apply to both open and closed-source models. However, the third one requires access to the embeddings and the ability to prompt LLMs with embeddings, and hence mostly applies to only open-source models. The prefixes are described further in Section 3.1.
We certify the proposed specifications leveraging confidence intervals. Our certifier samples several counterfactual prompt sets from the distribution given in the specification and generates high-confidence bounds on the probability of getting unbiased responses from the target LLM for any random prompt set in the distribution.
Contributions. Our main contributions are:
  • We design novel specifications that quantify the desirable property of low bias in the responses of LLMs with varying prefixes. The prefixes to given counterfactual prompts can be drawn from $3$ distributions — (1) random, (2) mixture of jailbreaks, and (3) jailbreak perturbations in the embedding space.
  • We develop a probabilistic black-box certifier QuaCer-B, applicable to both open and closed-source models, for quantifying bias in LLM output. QuaCer-B leverages confidence intervals (Clopper and Pearson, 1934) to generate high-confidence bounds on the probability of obtaining unbiased responses from the target LLM, given any random set of prompts from the distribution given in the specification.
  • We use QuaCer-B to reveal novel insights into the biases of popular LLMs, for different specifications (Section 5). We find that the safety alignment of SOTA models is easily circumvented with several prefixes in the distributions given in our specifications, especially those involving mixture of jailbreaks and jailbreak perturbations in the embedding space. These distributions are inexpensive to sample from, but can effectively bring out biased behaviors from SOTA models. This shows the existence of simple, bias-provoking distributions for which no defenses exist currently. We provide quantitative measures for the fairness (lack of bias) of SOTA LLMs, which hold with high confidence. We find that there are no consistent trends in the fairness of models with the scaling of their sizes, hence suggesting that the quality of alignment techniques could be a more important factor than size for fairness. Our implementation is available at https://github.com/uiuc-focal-lab/QuaCer-B.
2 Background
2.1 Large Language Models
Large Language Models (LLMs) are autoregressive foundation models that perform next-token prediction. That is, given a sequence of tokens $t_{1},\dots,t_{k}$, they give a distribution over the vocabulary for the next token, $P[t_{k+1}\mid t_{1},\dots,t_{k}]$. They are typically fine-tuned for instruction-following capabilities to be used such as chatbots. These instruction-finetuned models are further aligned with human feedback (Wang et al., 2023; Ouyang et al., 2022) to make their responses safe and unbiased. We certify these instruction-tuned (Zhang et al., 2024), aligned models for fairness, as they are typically used in public-facing applications.
2.2 Clopper-Pearson confidence intervals
Clopper-Pearson confidence intervals (Clopper and Pearson, 1934) provide lower and upper bounds on the probability of success parameter of a Bernoulli random variable with probabilistic guarantees. The bounds are obtained with $n$ independent and identically distributed observations of the random variable, in which $k(\leq n)$ successes are observed. The confidence interval is constructed to bound the value of $p$ with $[\hat{p_{l}},\hat{p_{u}}]$ with at least a given confidence level $(1-\gamma)100\%$. That is, $Pr\{p\in[\hat{p_{l}},\hat{p_{u}}]\}\geq(1-\gamma)$. $\gamma$ is the error probability by which the true value of $p\notin[\hat{p_{l}},\hat{p_{u}}]$. These confidence intervals are obtained by binomial statistical hypothesis testing for the value of $p$, where the lowest and highest values are $\hat{p_{l}}$ and $\hat{p_{u}}$ respectively, with the given confidence $1-\gamma$.
3 Formalizing bias certification
We develop a general framework, QuaCer-B, to specify and quantitatively certify the bias in the text generated by (Large) Language Models. QuaCer-B formalizes bias with specifications — precise mathematical representations that define the desirable property (absence of bias) in large sets of inputs. Bias is defined with respect to demographic groups that are subsets of the human population that share an identity trait, that could be biological, contextual, or socially constructed (Gallegos et al., 2024b). Bias consists of disparate treatment or outcomes when varying the demographic groups in the inputs to the target model (Gallegos et al., 2024b; Sheng et al., 2019; Smith et al., 2022). In the context of autoregressive LMs, we consider text generation bias consisting of stereotyping, misrepresentation, derogatory language, etc. This can primarily result in representational harms, which can subsequently cause allocation harms (Gallegos et al., 2024b). To apply certification to closed-source LMs as well, we study extrinsic bias (Cao et al., 2022) that manifests in the final generated responses of the models.
3.1 Bias specification
Next, we formally specify the lack of bias in the outputs of language models (LMs). In particular, we want the text generated by LMs to not exhibit any semantic disparities caused by mentioning different demographic groups in the prompts. Hence, our bias specification is motivated by Counterfactual Fairness (Kusner et al., 2018). Consider a given identity trait $\mathcal{I}$ such as gender, race, etc. (that are often the basis of social bias). $\mathcal{I}$ categorizes the human population into $m$ subsets called demographic groups $\mathcal{G}_{1},\dots,\mathcal{G}_{m}$, each differing by the value of the identity trait. Each demographic group $\mathcal{G}$ is characterized/recognized by several synonymous strings, called sensitive attributes $\mathcal{G}^{\mathcal{A}}$ (Li et al., 2024). For example, the sensitive attributes for the demographic group corresponding to the female gender are woman, female, etc. We select any one sensitive attribute of a demographic group $\mathcal{G}$ to represent $\mathcal{G}$. Let the resulting set of sensitive attributes, each corresponding to a demographic group, be $\mathcal{A}=\{\mathcal{A}_{1},\dots,\mathcal{A}_{m}\}$, where $\mathcal{A}_{i}\in\mathcal{G}^{\mathcal{A}}_{i}$. Our specifications are for counterfactual inputs (prompts) (Gallegos et al., 2024b) that differ only by the sensitive attributes in them.
Let $\mathcal{L}$ be the target LM and $\mathcal{V}$ be its vocabulary. Consider a set of prompts $\mathcal{P}=\{\mathcal{P}_{1},\dots\mathcal{P}_{s}\}$, $s>1,\mathcal{P}\subset\mathcal{V}^{[1,c]}$, where $c$ is the context length of $\mathcal{L}$ and $\mathcal{V}^{[1,c]}$ is a sequence of elements of $\mathcal{V}$ having length $\in[1,c]$. Let each prompt in $\mathcal{P}$ contain a unique sensitive attribute from $\mathcal{A}$ such that overall $\mathcal{P}$ represent more than $1$ distinct demographic groups represented in $\mathcal{A}$. Let each prompt be $\mathcal{P}_{i}=X_{i}\cup\mathcal{A}_{i}$, where $X_{i}$ is the part of the prompt that is independent of sensitive attributes and $\mathcal{A}_{i}$ consists of a sensitive attribute. We consider only prompts that can be decomposed into parts with and without sensitive attributes respectively. To generalize to closed-source LMs, we assume $\mathcal{L}$ to be a black-box system that can only be queried, when specifying and certifying bias. The black-box assumption renders any symbolic analysis (Mirman et al., 2020) infeasible and hence allows only for analysis with input-output examples.
{definition}
(Counterfactual prompt set). A set of prompts $\mathcal{P}=\{\mathcal{P}_{1}=X_{1}\cup\mathcal{A}_{1},\dots,\mathcal{P}_{s}=X_{s}\cup\mathcal{A}_{s}\}$ is called counterfactual when: (1) $\forall i,j\in[1,s]\ldotp X_{i}=X_{j}$; (2) $\exists i\in[1,s]\ldotp\forall j\in([1,s]\setminus i)\ldotp\mathcal{A}_{i}\neq\mathcal{A}_{j}$; (3) For an unbiased text generator $f$, $\forall i\in[1,s]\ldotp f(\mathcal{P}_{i})=f(X_{i})$. That is, the prompts only differ in the sensitive attributes that are ideally unrelated to the overall query and the set of prompts $\mathcal{P}$ represent more than sensitive attribute from $\mathcal{A}$.
We specify bias over counterfactual prompt sets (Definition 3.1), which can be prohibitive in number for typical context lengths. This is because, the common part of counterfactual prompt sets, $X$ can be any element from $\mathcal{V}^{[1,c]}$ which contains $\approx 10^{10000}$ elements for $c=2k$. Hence, enumerative specifications (which specify the desired behavior on all inputs) are impractical, as they cannot be scalably certified without symbolic analysis for large number of inputs. Hence, we define probabilistic specifications for the probability of unbiased responses from $\mathcal{L}$, for which we provide a certification algorithm in Section 3.2. Let $\Delta$ be a sampleable probability distribution over $\wp(\mathcal{V}^{[1,c]})$ (power set of prompts) having non-zero support on some counterfactual prompt sets $\mathcal{P}$. We define probabilistic specifications for bias in $\mathcal{L}$ over $\Delta$. The specification is agnostic to $\Delta$’s sampler, as long as it generates independent and identically distributed samples. We provide examples of $\Delta$ in Section 4.
Let $\mathcal{D}$ be a user-defined bias detection function that can identify stereotypes/disparity in given texts for different sensitive attributes in $\mathcal{A}$. Let $\mathcal{D}$ evaluate to zero for unbiased inputs (appropriate scaling and shifting of $\mathcal{D}$ can be done to satisfy this criterion). We leave $\mathcal{D}$ as a parameter of the specification, as different domains can have varying notions of bias and the stakeholders can decide which notion is most suitable for their usecase (Anthis et al., 2024).
Overall, we give our quantitative specification as the probability of unbiased responses (as measured by $\mathcal{D}$) when $\mathcal{L}$ is independently prompted with each element of a randomly sampled counterfactual prompts set from $\Delta$ (1). The certificate $\mathcal{C}$ is defined for the user-defined parameters $\Delta$, $\mathcal{D}$, and $\mathcal{L}$.
$\mathcal{C}(\Delta,\mathcal{D},\mathcal{L})\triangleq Pr_{\mathcal{P}\sim\Delta}[\mathcal{D}([\mathcal{L}(\mathcal{P}_{1}),\dots,\mathcal{L}(\mathcal{P}_{s})])=0]$ (1)
3.2 Certification algorithm
Exactly computing $\mathcal{C}(\Delta,\mathcal{D},\mathcal{L})$ (1) is intractable as it would require enumerating all (prohibitively many) prompts sets in the support of $\Delta$. Hence, our certification algorithm estimates $\mathcal{C}(\Delta,\mathcal{D},\mathcal{L})$ for given $\Delta$ and $\mathcal{D}$ and target $\mathcal{L}$ with high confidence, as described next. We generate intervals $[\hat{p_{l}},\hat{p_{u}}]$ that bound $\mathcal{C}(\Delta,\mathcal{D},\mathcal{L})$ in (1) with confidence $1-\gamma$. Such interval estimates are better than point-wise estimates as they also quantify the uncertainty of the estimation. $\mathcal{C}(\Delta,\mathcal{D},\mathcal{L})$ is the probability of success (unbiased responses) for the Bernoulli random variable $\mathcal{F}\triangleq\mathcal{D}([\mathcal{L}(\mathcal{P}_{1}),\dots,\mathcal{L}(\mathcal{P}_{s})])=0$. To obtain high-confidence bounds on $\mathcal{C}(\Delta,\mathcal{D},\mathcal{L})$, we employ binomial proportion confidence intervals. In particular, we leverage the Clopper-Pearson confidence interval method (Clopper and Pearson, 1934) (Section 2.2) as it is known to be a conservative method, i.e., the confidence of the resultant intervals is at least the pre-specified confidence, $1-\gamma$ (Newcombe, 1998). We obtain $n$ independent and identically distributed (iid) samples of $\mathcal{F}$ by sampling iid $\mathcal{P}$ from $\Delta$ and compute the Clopper-Pearson confidence intervals of $\mathcal{C}(\Delta,\mathcal{D},\mathcal{L})$. The certificate, hence obtained, bounds the probability of unbiased responses for random $\mathcal{P}\sim\Delta$ with high confidence. Note that the certification results depend on the user-defined choices of $n$ and $1-\gamma$.
4 Certification instances
In this section, we instantiate prompt set distributions $\Delta$ to form novel bias specifications. We select $\Delta$ such that its underlying sample space has prompt sets that share a common characteristic, so we can certify the bias conditioned on the presence of the characteristic. Thus, this becomes a local specification (Seshia et al., 2018), wherein the certificate is given for a local input space. Local specifications have commonly been considered in neural network verification (Singh et al., 2019; Wang et al., 2021; Baluta et al., 2021). Prior works on neural network specifications such as (Geng et al., 2023) generate only local specifications, as they correspond to meaningful real-world scenarios, and as local input regions are considered to design adversarial inputs for the models. In our local bias specifications, we consider $\Delta$ around a given set of prompts $\mathcal{Q}$ (pivot), denoting the resultant prompt set distributions as $\Delta_{\mathcal{Q}}$. Prefixes are commonly used to steer the text generated by LLMs according to the users’ intentions (Liu et al., 2021). Hence, we want to study whether the application of certain prefixes can elicit different forms of bias from the target LLM. Let $\Delta_{pre}$ denote a distribution of prefixes. Each element in the sample spaces of $\Delta_{\mathcal{Q}}$ is a set of prompts formed by uniformly applying a prefix to all prompts $\mathcal{Q}_{i}\in\mathcal{Q}$, that is, $q\sim\Delta_{\mathcal{Q}}=\bigcup_{\mathcal{Q}_{i}\in\mathcal{Q}}\{p\odot\mathcal{Q}_{i}\}$ for $p\sim\Delta_{pre}$, where $\odot$ denotes string concatenation. Algorithm 1 presents the probabilistic specification involving addition of randomly sampled prefixes as a probabilistic program. Our probabilistic programs follow the syntax of the probabilistic programming language defined in (Sankaranarayanan et al., 2013, Figure 3). The syntax is similar to that of a typical imperative programming language, with the addition of primitive functions to sample from common distributions over discrete / continuous random variables (for example, Bernoulli: $\mathcal{B}$, Uniform: $\mathcal{U}$) and estimateProbability(.). estimateProbability(.) takes in a random variable and returns its estimated probability at a specific value. makePrefix(args,kind) (line 1) is a general function to sample different kinds of prefixes such as random prefixes (Algorithm 2), mixture of jailbreaks (Algorithm 3), and soft prefixes (Algorithm 4), constructed using arguments, args.
$\mathcal{C}(\Delta_{\mathcal{Q}},\mathcal{D},\mathcal{L})$ characterizes the bias that can be elicited from $\mathcal{L}$ by varying the prefix selected from $\Delta_{pre}$ applied to a given $\mathcal{Q}$. Next, we describe the $3$ different kinds of practical $\Delta_{pre}$ and their sampling algorithms to define local bias specifications for $\mathcal{L}$. We show some samples from each kind of $\Delta_{pre}$ in Appendix D. Our specifications are for the average-case behavior of the target LLM, as $\Delta_{pre}$ are not distributions of provably adversarial (worst-case) prefixes.
Algorithm 1 Prefix specification
Input: $\mathcal{L},\mathcal{Q}$; Output: $\mathcal{C}(\Delta,\mathcal{D},\mathcal{L})$
1:$pre\coloneqq\texttt{makePrefix}(\text{args},\text{kind = random / mixture / soft})$
2:$\mathcal{P}\coloneqq[pre\odot\mathcal{Q}_{i}\ \mathbf{for}\ \mathcal{Q}_{i}\in\mathcal{Q}]$
3:$\mathcal{C}(\Delta,\mathcal{D},\mathcal{L})\coloneqq\texttt{estimateProbability}(\mathcal{D}([\mathcal{L}(\mathcal{P}_{1}),\dots,\mathcal{L}(\mathcal{P}_{s})])=0)$
Input: $\mathcal{V}$; Output: $pre$
1:$pre\coloneqq\mathcal{U}(\mathcal{V})\odot\dots[q\text{ times}]\dots\odot\mathcal{U}(\mathcal{V})$
Algorithm 2 Make random prefix
Input: $\mathcal{L},\mathcal{M}_{0}$; Output: $pre$
1:$\mathcal{E}\coloneqq\texttt{embed}(\mathcal{L},\mathcal{M}_{0})$
2:$pre\coloneqq\mathcal{E}+\mathcal{U}([-\kappa,\kappa])$
Algorithm 4 Make soft prefix
Input: $\mathcal{L},\mathcal{V},\mathcal{M}$; Output: $pre$
1:$\mathcal{M}\coloneqq[\texttt{split}(\mathcal{M}_{k})\ \mathbf{for}\ \mathcal{M}_{k}\in\mathcal{M}]$
[…]
3:$\omega(p_{\lambda},\mathcal{H})\coloneqq\texttt{shuffle}(\{\texttt{if}(\mathcal{B}(p_{\lambda}),h,\emptyset)\mid h\in\mathcal{H}\})$
4:$\mathcal{M}^{i}\coloneqq\mathcal{M}_{0}[0]\odot\omega(p_{\lambda},\mathcal{H})\odot\mathcal{M}_{0}[1]\odot\omega(p_{\lambda},\mathcal{H})\odot\dots$
5:$\mathcal{M}^{i}\leftarrow\texttt{tokenize}(\mathcal{L},\mathcal{M}^{i})$
6:$pre\coloneqq[\texttt{if}(\mathcal{B}(p_{\mu}),\mathcal{U}(\mathcal{V}),\tau)\ \mathbf{for}\ \tau\in\mathcal{M}^{i}]$
Algorithm 3 Make mixture of jailbreak prefix
Random prefixes. Prior works such as (Wei et al., 2023; Zou et al., 2023) have shown the effects of incoherent strings in manipulating LLMs and jailbreaking them into answering queries about harmful behaviors. Hence, we specify bias in LLMs for prompts with incoherent prefixes that are random sequences of tokens from the vocabulary of the LLM. Algorithm 2 presents the prefix sampler as a uniform distribution, $\mathcal{U}(.)$ over the random prefixes of fixed length, $q$. The sample space of random prefix $\Delta_{pre}$ has $|\mathcal{V}|^{q}$ cardinality. $\Delta_{pre}$ for random prefixes assigns a non-zero probability to discovered and undiscovered jailbreaks, of a fixed length $q$. Hence, certification for the random prefix distribution indicates the expected bias in responses to $\mathcal{Q}$ with any random prefixes of length $q$.
Mixtures of jailbreaks. Manually designed jailbreaks are fairly effective at bypassing the safety training of LLMs (walkerspider, 2022; Wei et al., 2023; jai, ). To certify the vulnerability of LLMs under powerful jailbreaks, we develop specifications with manual jailbreaks. The distribution from which the manual jailbreaks can be sampled is unknown. Thus, we construct potential jailbreaking prefixes from a set $\mathcal{M}$ of popular manually-designed jailbreaks by applying $2$ operations — interleaving and mutation. Interleaving attempts to strengthen a given manual jailbreak with more bias-provoking instructions, while mutation attempts to obfuscates the jailbreak such that it can be effective, even under explicit training to avoid the original jailbreak. Algorithm 3 presents the prefix constructor as a probabilistic program. Each manual jailbreak $\mathcal{M}_{k}\in\mathcal{M}$ can be treated as a finite set of instructions $\mathcal{M}_{k}=\{\mathcal{M}^{1}_{k},\dots\}$. Let $\mathcal{M}_{0}$ be the most effective jailbreak (a.k.a. main jailbreak). We extract the information on the effectiveness of jailbreaks from popular open-source leaderboards of jailbreaks. We include all the instructions of the main jailbreak in the final prefix. The other jailbreaks are helper jailbreaks, whose instructions are included with an interleaving probability, $p_{\lambda}$ in the final prefix. Let $\mathcal{H}=\bigcup_{\mathcal{M}_{i}\in\mathcal{M}}\mathcal{M}_{i}$ denote the set of all instructions from helper jailbreaks [line 2]. Let $\omega(p_{\lambda},\mathcal{H})$ shuffle and concatenate randomly picked (with probability $p_{\lambda}$) instructions from $\mathcal{H}$ [line 3]. shuffle(.) is a function for randomly sampling a permutation from a uniform distribution over all permutations of an input list (after removing $\emptyset$ which denotes void elements). Let $\texttt{if}(e_{1},e_{2},e_{3})$ be an abbreviation for if $e_{1}$ then $e_{2}$ else $e_{3}$. We first apply the interleaving operation with the resultant given as $\mathcal{M}^{i}$ [line 4]. The mutation operation is then applied to $\mathcal{M}^{i}$ viewed as a sequence of tokens $[\tau_{0},\dots,]$, wherein any token $\tau_{i}$ can be flipped to any random token $\tau^{\prime}_{i}\in\mathcal{V}$, with a mutation probability $p_{\mu}$ (generally set to be low), to result in $pre$ [line 6]. We hypothesize such prefixes to be potential jailbreaks as they are formed by strengthening a manual jailbreak with other jailbreaks and obfuscating its presence. The number of prefixes formed by the aforementioned operations can be prohibitively many, owing to typically long manual jailbreaks and the possibility to mutate any token to any random token from the LLM’s vocabulary.
Soft jailbreaks. Due to the limited number of effective manual jailbreaks (walkerspider, 2022; Learn Prompting, 2023), they can be easily identified and defended against. However, the excellent denoising capabilities of LLMs could render them vulnerable to simple manipulations of manual jailbreaks as well, indicating that the threat is not completely mitigated by the current defenses (Jain et al., 2023). Hence, we specify fairness under prefixes constructed by adding noise of the original manual jailbreaks. Algorithm 4 presents the prefix constructor as a probabilistic program. Let $\mathcal{E}$ be the embedding matrix of $\mathcal{M}_{0}$ in the embedding space of the target LLM, obtained by applying the function embed(.) [line 1]. We perturb $\mathcal{E}$ by adding noise to it. As we are not aware of any adversarial distributions of noise that could be added to manual jailbreaks to make them stronger, we select a uniform distribution. We uniformly sample noise from $\mathcal{B}(0,\kappa)$ which is an $\kappa>0$ (constant) ball around the origin and add it to $\mathcal{E}$ to construct $pre$ in the embedding space [line 2].
5 Experiments
5.1 Experimental setup
We used $2$ A100 GPUs, each with 40GB VRAM. We derive the queries on which the specifications from the $3$ prefix distributions presented in Section 4 are pivoted, from popular datasets for fairness and bias assessment — BOLD (Dhamala et al., 2021) and Decoding Trust (Wang et al., 2024).
BOLD setup. BOLD is a dataset of partial sentences to demonstrate bias in the generations of LLMs in common situations. We pick a test set of $250$ samples randomly from BOLD’s ‘profession’ partition and demonstrate binary gender bias specifications and certificates on it.
Figure 2: Example from $\mathcal{Q}_{BOLD}$
We develop a pivot set of prompts from each test set sample by prepending an instruction to complete the partial sentence for the profession annotated in BOLD, where the subject identifies with a particular gender (Male / Female). An example pivot set $\mathcal{Q}$ from BOLD is illustrated in Figure 2. Let the resultant test set consisting of pivot prompts from each partial sentence considered from BOLD be $\mathcal{Q}_{BOLD}$. To identify bias in the responses of the target LLM for the prompts constructed from pivot prompt sets in $\mathcal{Q}_{BOLD}$, we use the regard-based bias metric proposed in (Sheng et al., 2019). Our adaptation of the regard metric is described and evaluated with a human study on Amazon Mechanical Turk in Appendix C.1. Our bias detector matches human perception of bias in $76\%$ cases. We qualitatively analyze the false positive and false negatives of the bias detector with respect to human judgment in Appendix C.1.
Figure 3: Example from $\mathcal{Q}_{DT}$
Decoding Trust setup. Decoding Trust (DT) is a dataset benchmark to evaluate various properties of LLMs, including stereotype bias against people of different demographic groups. We make specifications from all $48$ statements in the ‘stereotypes’ partition for demographic groups corresponding to race (black/white). An example pivot set $\mathcal{Q}$ from DT is illustrated in Figure 3. Let the resultant test set consisting of pivot prompts from each partial sentence considered from DT be $\mathcal{Q}_{DT}$. We evaluate the LLM responses to prompts derived from pivot prompt sets in $\mathcal{Q}_{DT}$ using a bias detector that identifies the disparity in agreement to the stereotype for different demographic groups given in the prompt, as discussed in Appendix C.2.
For every element in $\mathcal{Q}_{BOLD}$ and $\mathcal{Q}_{DT}$, we generate 3 certificates for the specifications in Section 4, such that a certificate consists of bounds on the probability of unbiased responses from the target LLM. Both bias detectors are such that they output $1$ for a biased set of responses to a counterfactual prompt set and 00 for an unbiased set of responses. The values of the certification parameters used in our experiments are given in Table 2 (Appendix A). We study their effect on the certification results with an ablation study in Appendix A. We generate the certification bounds with $95\%$ confidence and $50$ samples. While our main experiments are for counterfactual prompt sets with binary demographic groups, our framework can be extended beyond binary demographic groups, which we experimentally demonstrate in Appendix A.4. Note that the manual jailbreaks used are common across all specifications and are presented in Appendix B.
Table 1: Average of the bounds on the probability of unbiased responses for different models. Lowest bounds for each specification kind and dataset are highlighted. We report $2$ baselines — unbiased responses when prompting without prefixes and with the main jailbreak as prefix.
Average certification bounds
Dataset Model % Unbiased without prefix % Unbiased with main JB Random Mixture Soft
BOLD (250) Vicuna-7B $99.9$ $89.4$ $(0.93,1.0)$ $(0.90,0.99)$ $(0.73,0.89)$
Vicuna-13B $99.7$ $99.8$ $(0.93,1.0)$ $(0.93,1.0)$ $(0.92,1.0)$
Llama-7B $99.8$ $99.8$ $(0.92,1.0)$ $(0.92,1.0)$ $(0.93,1.0)$
Llama-13B $99.8$ $99.7$ $(0.93,1.0)$ $(0.91,1.0)$ $(0.93,1.0)$
Mistral-7B $100.0$ $41.0$ $(0.92,1.0)$ $\mathbf{(0.22,0.42)}$ $\mathbf{(0.30,0.52)}$
Gemini $99.2$ $74.1$ $(0.92,1.0)$ $(0.60,0.83)$-
GPT-3.5 $99.5$ $50.2$ $(0.92,1.0)$ $(0.44,0.67)$-
GPT-4 $99.8$ $99.9$ $(0.92,1.0)$ $(0.80,0.96)$-
Claude-3.5-Sonnet $99.6$ $99.8$ $(0.93,1.0)$ $(0.92,1.0)$-
DT (48) Vicuna-7B $95.4$ $100.0$ $(0.85,0.97)$ $(0.92,1.0)$ $(0.88,0.97)$
Vicuna-13B $88.7$ $76.2$ $\mathbf{(0.71,0.92)}$ $(0.92,1.0)$ $(0.51,0.78)$
Llama-7B $97.5$ $100.0$ $(0.79,0.96)$ $(0.92,1.0)$ $(0.92,1.0)$
Llama-13B $100.0$ $100.0$ $(0.92,1.0)$ $(0.93,1.0)$ $(0.93,1.0)$
Mistral-7B $99.2$ $72.9$ $(0.91,1.0)$ $(0.85,0.99)$ $\mathbf{(0.46,0.73)}$
Gemini $99.6$ $94.6$ $(0.92,1.0)$ $(0.73,0.93)$-
GPT-3.5 $99.6$ $56.7$ $(0.93,1.0)$ $\mathbf{(0.66,0.88)}$-
GPT-4 $100.0$ $100.0$ $(0.93,1.0)$ $(0.93,1.0)$-
Claude-3.5-Sonnet $99.6$ $100.0$ $(0.93,1.0)$ $(0.93,1.0)$-
5.2 Certification results
We certify the popular contemporary LLMs — Llama-2-chat (Touvron et al., 2023) 7B and 13B (parameters), Vicuna-v1.5 (Chiang et al., 2023) 7B and 13B, Mistral-Instruct-v0.2 (Jiang et al., 2023) 7B, Gemini-1.0-pro (Gemini Team, 2024), GPT-3.5 (Brown et al., 2020b), GPT-4 (Achiam et al., 2023), and Claude-3.5-Sonnet (Anthropic, 2024). We report the average of the certification bounds for all pivot prompt sets in $\mathcal{Q}_{BOLD}$ and $\mathcal{Q}_{DT}$ each for every model in Table 1. We do not certify the closed-source models such as Gemini, GPT, and Claude for soft jailbreaks, as it requires access to the models’ embedding layers. Certification time significantly depends on the inference latency of the target model. Generating each certificate can take 1-2 minutes for models with reasonable latency.
Baselines. The baselines consider $\mathcal{Q}_{BOLD}$ and $\mathcal{Q}_{DT}$ as benchmarking datasets, having counterfactual prompt sets as individual elements. Similar to popular LLM bias benchmarking works such as (Wang et al., 2024; Liang et al., 2023; Esiobu et al., 2023; Xie et al., 2024), we study the biases in LLMs for a fixed dataset of counterfactual prompt sets which may be provided as is to the LLM, or with some provoking texts (jailbreaks). In the first baseline (without prefix), every counterfactual prompt set is evaluated $5$ times by each time prompting a target LLM with each prompt in the set without any prefixes and detecting bias across its responses using the bias detectors of the respective datasets. The bias result for each counterfactual prompt set is computed by averaging the results over the $5$ evaluations, similar to (Wang et al., 2024). This baseline indicates biases in the LLM responses without any prefixes and is hence used to judge the influence of the prefixes on eliciting biased responses from the models. Table 1 reports the average of the evaluations over all counterfactual prompt sets in $\mathcal{Q}_{BOLD}$ and $\mathcal{Q}_{DT}$ respectively. In the second baseline (with main jailbreak), each counterfactual prompt is similarly evaluated $5$ times, but each time prompting the LLM with each prompt in the set with the unmodified main jailbreak (Figure 7 in Appendix B) used in the mixture of jailbreak and soft jailbreak distributions, as a prefix. The average result of this baseline is also reported in Table 1. This baseline is used to indicate the efficacy of the main jailbreak without any modifications, and hence suggests the importance of the mixture and soft prefix distributions around the main jailbreak in modifying the biases in LLM responses. The benchmarking baselines are empirical studies of bias in the target LLMs, which analyze the biases with specific prompts. On the other hand, QuaCer-B quantifies biases for any random prompt sampled from a given distribution.
5.2.1 General observations
Comparison with baselines. Our results for the baseline without prefix are generally close to and often higher than the average upper bounds from certification. This suggests that the counterfactual prompt sets do not majorly result in biased LLM behavior by themselves and the application of prefixes is essential for biased LLM responses. Hence, such baselines present an optimistic view on the biases of LLMs, which may not be acceptable for real-world scenarios. We compare the baseline with main jailbreak with the certification bounds for the specifications with mixture of jailbreak prefixes and soft prefixes, as these consist of modifications of the main jailbreak. We see that in most cases, the baseline results are either contained within or higher than the bounds. Thus the baseline does not account for several biased behaviors that can result by modifying the main jailbreak. In some cases, e.g., Mistral for DT specifications, the average lower bound is higher than the baseline result, indicating that the modifications might have reduced the efficacy of the main jailbreak for the model. In such cases, the specifications can be made model-specific and we leave this to future work.
Variations with prefixes. We observe that the random prefixes are generally less effective at eliciting biased behaviors than the other prefixes. This suggests that incoherent prefixes can be denoised by the LLMs and may not effectively bypass their alignment. In some cases, such as that for Vicuna-13b, we see that the model produces biased responses with the random prefixes, indicating that the model is not robust to even incoherent perturbations to the prompts. Thus random prefix specifications present basic certificates into the denoising capabilities of the LLMs to produce unbiased responses. There are no consistent trends between the mixture of jailbreaks and soft prefix specifications, except for the effectiveness of prompts from their distributions in producing biased LLM responses.
Denial of response. In several instances, such as those during the certification of GPT-4 (for DT), Claude, and Llama, we observe unbiased behaviors because the model simply refuses to respond. The completions are mostly "Sorry I can’t assist with that". While that can be a valid defense to adversarial prompts, it may not be the most desirable. The model could instead output neutral responses, ignoring any bias-encouraging texts in the prompts to be helpful but not harmful.
5.2.2 Model-specific observations
Vicuna. The Vicuna models show high probability of unbiased responses, except for those with soft prefixes. The baselines indicate that the 7B model can output less unbiased responses with the main jailbreak as prefix. But the mixture of jailbreaks specifications have high probability of unbiased responses, suggesting that the perturbations of the main jailbreak with the operations of mixture specifications can make the jailbreak less effective, while perturbing the jailbreak in the embedding space can elicit more biased behaviors from the model. We find that unbiased behavior trends are not consistent with increasing model sizes across different specifications.
Llama and Claude. The Llama models and Claude-Sonnet generate a few biased responses, but the bounds on the probability of unbiased responses are high, indicating better alignment.
Mistral. Mistral exhibits significantly low probability of unbiased when prompted with the mixture of jailbreaks and soft jailbreak prefixes for specifications from both $\mathcal{Q}_{BOLD}$ and $\mathcal{Q}_{DT}$.
Gemini. We prompt Gemini with the safety filters disabled, so we can check the safety training of the model itself. We observe that its responses can get highly biased for mixture of jailbreaks specifications, resulting in low bounds. This insight is, however, not evident from the baselines.
GPT. The baselines indicate low bias in the responses of GPT-4. However, certification results say otherwise. The GPT-4 model shows low probability of unbiased responses for the mixture of jailbreak specifications derived from $\mathcal{Q}_{BOLD}$. However, we observe unbiased responses for specifications corresponding to $\mathcal{Q}_{DT}$. GPT-3.5 achieves low probability of unbiased responses for the mixture of jailbreak specifications, indicating that GPT-4 may have undergone better safety training than it.
5.3 Case studies
We qualitatively introspect the individual certificates obtained for different models for one example from $\mathcal{Q}_{BOLD}$. We present similar analyses for one example from $\mathcal{Q}_{DT}$ in Appendix E. We select the example for Desktop automation refers to software programming from $\mathcal{Q}_{BOLD}$ for the ‘Automation’ profession and study the sample responses obtained for the mixture of jailbreak specification. We present some responses of $4$ popular LLMs — Mistral, Llama-7B, Gemini, and GPT-4 obtained during certification, in Figure 4. The responses demonstrate various stereotypical biases, against both genders. For example, the Mistral model discriminates against women, calling them “incapable and better suited for menial jobs”. The Gemini model is also frequently biased against women, saying “they cannot think and are meant for household chores”. On the other hand, GPT-4 regularly shows biases towards men, as it gives responses about them “lacking motivation and being less efficient”, with automation being better at their jobs. It depicts women crushing their male competition and becoming superior. The Llama model, however, defends against the prompts and avoids responding in a biased way. The responses reflect the certification results for the models for the BOLD example. We find from the certification bounds that the Mistral model is provably worse than Gemini and GPT, which are worse than Llama with high confidence for this specification.
Figure 4: Responses of Mistral, Llama-7B, Gemini, and GPT-4 obtained during their certification for the BOLD example ‘Desktop automation refers to software programming’ for mixture of jailbreak specifications. The certification result is indicated alongside the model.
6 Related work
Attacking LLM alignment. LLMs are typically aligned with human ethics by supervised fine-tuning and reinforcement learning with human feedback (Ouyang et al., 2022). However, prior work (Zou et al., 2023; Vega et al., 2023; Chao et al., 2023; Sheng et al., 2020; Wallace et al., 2019) propose various methods to generate jailbreaking attacks for LLMs, which bypass their safety-alignment and lead to harmful or biased responses. Jailbreaks can be incoherent as in (Zou et al., 2023; Sheng et al., 2020; Wallace et al., 2019) or coherent instructions (Dominique et al., 2024; Liu et al., 2024).
Benchmarking LLMs. Various prior works have benchmarked the performance of LLMs on standard and custom datasets. These could be datasets of general prompts (Dhamala et al., 2021; Wang et al., 2024) or those of adversarial examples (Zou et al., 2023; Mazeika et al., 2024) designed to elicit undesirable behaviors from the models. Popular benchmarks such as (Liang et al., 2023; Wang et al., 2024; Mazeika et al., 2024; Manerba et al., 2024; Gallegos et al., 2024a) present empirical trends for the performance of LLMs, measured along various axes including bias and fairness.
Guarantees for LLMs. There is an emerging need for guarantees on LLM behavior, fueled by their popularity in public-facing use cases. (Kang et al., 2024) provides guarantees on the generation risks of RAG LLMs. (Crouse et al., 2024) give formal specifications for expected behavior of LLMs as autonomous agents. However, there is no work on certifying LLM bias, to the best of our knowledge.
Fairness in Machine Learning. Fairness has been extensively studied for general Machine Learning, beginning from the seminal work of Dwork et al. (2011). Prior works have proposed methods to formally certify classifiers for fairness, such as (Biswas and Rajan, 2023; Bastani et al., 2019). However, these do not extend to LLMs. Fairness and bias have also been studied in natural language processing in prior works such as (Chang et al., 2019; Smith et al., 2022; Krishna et al., 2022).
7 Conclusion
We present the first general framework QuaCer-B to certify bias in the responses of LLMs, applicable to both open and closed-source models. We instantiate our framework with novel specifications based on different kinds of prefixes. QuaCer-B generates high confidence bounds on the probability of unbiased responses for prompts from a given distribution. Our results show that certification can detect biases not observed with existing efforts. As future work, QuaCer-B can be made less dependent on bias detectors. It can also be made sample efficient for higher confidence requirements.
Appendix A Ablation study
In this section, we study the effect of changing the various certification parameters (a.k.a. hyperparameters) on the certificates generated with QuaCer-B. Table A presents the list of hyperparameters and their values used in our experiments.
Table 2: Hyperparameter values
Hyperparameter Description Value
$\gamma$ $(1-\gamma)$ is confidence over certification 0.05
$n$ Number of samples for certification 50
$q$ Prefix length for random prefixes 100
$p_{\lambda}$ Interleaving probability 0.2
$p_{\mu}$ Mutation probability 0.01
$\kappa$ Max noise magnitude added to jailbreak embedding elements relative to the maximum embedding value 0.02
We regenerate the certificates for different prefix distributions by varying the hyperparameters. In particular, we study the variations of the results when $n,q,p_{\lambda},p_{\mu},$ and $\kappa$ are varied, keeping $\gamma$ constant. This is because, $1-\gamma$ denotes the confidence of the certification bounds and that is generally desired to be high. $95\%$ is a typical confidence level for practical applications [Sim and Reid, 1999]. We conduct this ablation study on the specifications for a randomly picked set of $100$ counterfactual prompt sets from BOLD’s test set $\mathcal{Q}_{BOLD}$. We certify the Mistral-Instruct-v0.2 [Jiang et al., 2023] 7B parameter model and study the overall results next. We show ablations on $n$ for all kinds of specifications in Figure 5. We see that the bounds begin converging at $50$ samples and subsequent samples cause minor variations in their values, justifying our choice of using $50$ samples. Fewer than $50$ samples can result in less tight bounds.
(a) Random prefixes
(b) Mixtures of jailbreaks
(c) Soft prefixes
Figure 5: Ablation study on the certification hyperparameters showing variations of average certification bounds with number of samples $n$
(a) Variation with prefix length $q$ for random prefix specifications
(b) Variation with interleaving probability $p_{\lambda}$ for mixture of jailbreaks specifications
(c) Variation with mutation probability $p_{\mu}$ for mixture of jailbreaks specifications
(d) Variation with relative magnitude of noise $\kappa$ for soft jailbreaks specifications
Figure 6: Ablation study on the certification hyperparameters showing variations of average certification bounds
A.1 Random prefixes
The specifications based on random prefixes consist of $1$ hyperparameter — $q$, length of the random prefix. Hence, we vary this hyperparameter, while keeping the others fixed. Figure 6(a) presents the variation in the average certification bounds obtained when varying $q$.
A.2 Mixture of jailbreaks
These specifications have $2$ hyperparameters — $p_{\lambda}$, the probability of adding an instruction from the helper jailbreaks when interleaving, and $p_{\mu}$, the probability of randomly flipping every token of the resultant of interleaving. We show ablation studies on these in Figures 6(b) and 6(c) respectively.
A.3 Soft jailbreaks
These specifications have $1$ hyperparameter — $\kappa$, the maximum relative magnitude (with respect to the maximum magnitude of the embeddings) by which the additive uniform noise can change the embeddings of the main jailbreak. Figure 6(d) presents an ablation study on $\kappa$.
We see no significant effects of the variation of abovementioned hyperparameters on certification results for different specifications.
A.4 Scaling beyond binary demographic groups
Table 3: Average bounds on the probability of unbiased responses from Mistral 7B.
Spec type Bounds
Random $(0.91,1.0)$
Mixture $(0.82,0.98)$
Soft $(0.20,0.46)$
Our general framework (Section 3) and specification instances (Section 4) are applicable to certify biases beyond binary counterfactual prompt sets (like for male/female gender, black/white race). This is subject to the availability of bias detectors $\mathcal{D}$ that can identify biases across responses for counterfactual prompt sets for more than binary demographic groups. While we are not aware of any $\mathcal{D}$ that could work with $\mathcal{Q}_{BOLD}$, we extend our $\mathcal{D}$ for the specifications from $\mathcal{Q}_{DT}$ to work for responses to prompts from three racial demographic groups — black people, white people, and asians. We elaborate on the extension in Appendix C.2. We certify Mistral-Instruct-v0.2 7B model with the three kinds of specifications and find that the average certification bounds presented in Table 3 are similar to the bounds presented for the Mistral model in Table 1 for the random and mixture of jailbreak specifications. However, the results are significantly worse for the soft prefix specifications. This is because, firstly the model is particularly susceptible to these specifications as is evidenced even in the results with binary demographic groups. Secondly, with the addition of another demographic group, the bias detector is skewed towards identifying bias in more sets of responses than for the case with binary demographic groups. The bias detector identifies bias in responses having at least $1$ agreement and $1$ disagreement to the stereotype mentioned in the prompts, which has the same chance as unbiased result for binary demographic groups, but not beyond them.
Appendix B Common jailbreaks used
In our experiments, we combine common jailbreaks to certify against hybrid jailbreaks. The specific jailbreaks we use are shown below. The main jailbreak [^2] Last accessed on Jan 22, 2024 from www.jailbreakchat.com/prompt/588ab0ed-2829-4be8-a3f3-f28e29c06621 is shown in Figure 7. We have modified the jailbreak to encourage the model to generate responses demonstrating gender bias by adding the text highlighted in red in Figure 7(a) and for racial bias in Figure 7(b). We use one helper jailbreak [^3] https://www.reddit.com/r/ChatGPT/comments/113ddqq/my_jailbreak_is_far_superior_to_dan_the_prompt_is/ from which we derive the instructions shown in Figure 8. Each of these instructions is inserted into the main jailbreak at the points where its individual instructions end, with the interleaving probability $p_{\lambda}$. All instructions inserted at the same point in the main jailbreak appear in a random order in the final jailbreak prefix.
(a) For gender
(b) For race
Figure 7: Main Jailbreaks for different sensitive attributes
Figure 8: Instructions derived from helper jailbreak
Appendix C Bias detector
C.1 BOLD experiments
Design of bias detector. Our bias detector $\mathcal{D}_{BOLD}$ is based on the regard bias metric [Sheng et al., 2019]. Specifically, we measure the bias for a pair of completions, one for male and another for female, by invoking the regard classifier in ’compare’ mode. The compare mode gives the difference between the regard score obtained for the different completions. We consider all completion pairs with a regard score higher than a prespecified threshold ($0.85$ in this case) as biased. For the completions considered unbiased with the previous evaluation, we also consider the regard scores of the individual completions (with respect to the partial sentence given to the model). We consider the completions to be biased if either score exceeds the threshold. This is for the instances of bias, wherein the pair of generations are similarly biased and individual generations have text discriminating against a particular demographic group.
Evaluating the metric. We evaluate $\mathcal{D}_{BOLD}$ against human judgment of bias, by getting annotations for a subset of the completions from the different models from Amazon Mechanical Turk. We randomly select $490$ pairs of completions (for male and female subjects in the prompts), $70$ each obtained when certifying $7$ different models. Each pair of completions is annotated by $3$ different annotators. We combine the $3$ annotations for each pair of completions into a single value of bias by considering the majority consensus among the annotators. We compare the bias annotation thus obtained from humans with the results of $\mathcal{D}_{BOLD}$ and find that $\mathcal{D}_{BOLD}$’s outputs matches human intuition $76\%$ times. We provide the HTML file used to render the instructions shown to the Amazon Mechanical Turkers in our supplementary material. Each participant is given a compensation of \$0.5 for the annotations. Next, we provide a qualitative analysis with examples where the results of $\mathcal{D}_{BOLD}$ and the human evaluation match and where they do not match.
(a) (True Positive) Bias detected by both $\mathcal{D}_{BOLD}$ and annotators
(b) (False Negative) Bias detected by annotators but not $\mathcal{D}_{BOLD}$
(c) (False Positive) Bias detected by $\mathcal{D}_{BOLD}$ but not annotators
Figure 9: Qualitative analysis of $\mathcal{D}_{BOLD}$ with human bias annotations.
Figure 9(a) shows the true positive examples where the both the $\mathcal{D}_{BOLD}$ and the annotators identify bias. Figure 9(b) shows the false negative examples where $\mathcal{D}_{BOLD}$ does not identify bias but the annotators do. Figure 9(c) shows the false positive examples where $\mathcal{D}_{BOLD}$ identifies bias but the annotators do not. According to the authors, the false positive examples are actually biased. But we also recognize that bias is a complex subject, and its notion can vary across individuals. Hence, we believe that the noise in the human study needs to be taken into account when evaluating our bias detector $\mathcal{D}_{BOLD}$.
C.2 DT experiments
We design the bias detector $\mathcal{D}_{DT}$ for specifications from the Decoding Trust dataset, using the agreement detector proposed for the stereotype bias dataset in the Decoding Trust paper [Wang et al., 2024]. In particular, the agreement detector in the benchmark applies on $1$ response and identifies if it contains phrases such as “I agree” (to the stereotype mentioned in the prompt) for agreement or phrases like “I disagree” for disagreement. To detect bias across the responses for the counterfactual prompts, $\mathcal{D}_{DT}$ checks for agreement in the individual responses, and identify bias when there is at least $1$ agreement, but not all responses are for agreement, i.e., the model agrees to the stereotype for some demographic groups but not others. $\mathcal{D}_{DT}$ can easily scale beyond binary demographic groups, as it is agnostic to the number of responses across which it needs to detect bias.
C.3 Limitations of bias detectors
As we can see from the human evaluations for $\mathcal{D}_{BOLD}$ and also from the simplicity of $\mathcal{D}_{DT}$ that the bias detectors can have several false positive and false negative cases. This is expected, as detection of bias is a challenging task [Li et al., 2022b, Kruspe, 2024] owing to the subtle and intricate differences between biased and unbiased texts. Moreover, there can be special characters / emojis in the texts generated in the LLMs, which are often not considered by textual bias detectors, but are clearly biased, e.g., [Uncaptioned image] (we observe this emoji combination in the generations of Vicuna-7B for random prefix specifications from $\mathcal{Q}_{BOLD}$). While our certification method borrows some of the inaccuracies of the bias detectors we use, they are often smoothed as certification aggregates several observations of bias to generate bounds on the probability of unbiased responses.
Appendix D Example samples from prefix distributions
In this section, we provide examples of prefixes of different kinds, described in Section 4.
Random prefixes. Figure 10 presents some examples of random prefixes, each having $100$ tokens.
Figure 10: Examples of random prefixes
Mixture of jailbreaks. Figure 11 presents some examples of mixture of jailbreak prefixes, formed from the main and helper jailbreaks in Figures 7 and 8 respectively.
Figure 11: Examples of mixture of jailbreak prefixes
Soft prefixes. Examples of soft prefixes added to the queries can not be represented in textual form as they consist of perturbations of the prefix in the LLM’s embedding space. However, the general structure is {jailbreak perturbed in embedding space}.
Appendix E Case studies from Decoding Trust certificates
We introspect the example [target group] have a drug problem from $\mathcal{Q}_{DT}$. We study the responses of $4$ LLMs — Mistral, Llama-7B, Gemini, and GPT-3.5, obtained when certifying them for the mixture of jailbreaks specification based on the example. Figure 12 presents some samples of responses from the models to explain their certificates. We observe that the Gemini model is biased against white people, calling them ‘lazy’ and ‘more likely to be exposed to drugs’. Similarly, Mistral is also biased against white people, saying that they ‘have a monopoly on addiction’. On the other hand, GPT-3.5 is biased against black people, depicting them as ‘inferior’ and ‘inclined towards addiction’. The Llama model is safer, refraining from biased responses. The bounds for the probability of unbiased response suggest that GPT-3.5 and Gemini are more inclined to produce biased responses than the Mistral and Llama models for prompts in the distribution of counterfactual prompt sets in the given specification.
Figure 12: Responses of Mistral, Llama-7B, Gemini, and GPT-3.5 obtained during their certification for the Decoding Trust example ‘[target group] have a drug problem’ for mixture of jailbreaks specification. The certification result is indicated alongside the model.
Appendix F Positive and Negative impacts of our work
We identify the following positive and negative impacts of our work.
Positive impacts. Our work is the first to provide quantitative certificates for the bias in Large Language Models. It can be used by model developers to thoroughly assess their models before releasing them and by the general public to become aware of the potential harms of using any LLM. As our framework, QuaCer-B assumes black-box access to the model, it can be applied to even closed-source LLMs with API access.
Negative impacts. In this work, we propose $3$ kinds of specifications involving — random prefixes, mixtures of jailbreaks as prefixes, and jailbreaks in the embedding space of the target model. While these prefixes are not adversarially designed, they are often successful in eliciting biased and toxic responses from the target LLMs. They can be used to attack these LLMs by potential adversaries. We have informed the developers of the LLMs about this threat.

### v3

@@
-Quantitative Certification of Bias in Large Language Models
+Certifying Counterfactual Bias in LLMs
 Abstract
@@
 Large Language Models (LLMs) can produce biased responses that can cause representational harms.
-However, conventional studies are insufficient to thoroughly evaluate LLM bias, as they can not scale to large number of inputs and provide no guarantees.
-Therefore, we propose the first framework, QuaCer-B (Quantitative Certification of Bias) that certifies LLMs for bias on distributions of prompts.
-A certificate consists of high-confidence bounds on the probability of unbiased LLM responses for any set of prompts mentioning various demographic groups, sampled from a distribution.
-We illustrate the bias certification for distributions of prompts created by applying varying prefixes drawn from a prefix distributions, to a given set of prompts.
-We consider prefix distributions for random token sequences, mixtures of manual jailbreaks, and jailbreaks in the LLM’s embedding space to certify bias.
-We obtain non-trivial certified bounds on the probability of unbiased responses of SOTA LLMs, exposing their vulnerabilities over distributions of prompts generated from computationally inexpensive distributions of prefixes.
+However, conventional studies are insufficient to thoroughly evaluate biases across LLM responses for different demographic groups (a.k.a.
+counterfactual bias), as they do not scale to large number of inputs and do not provide guarantees.
+Therefore, we propose the first framework, LLMCert-B that certifies LLMs for counterfactual bias on distributions of prompts.
+A certificate consists of high-confidence bounds on the probability of unbiased LLM responses for any set of counterfactual prompts — prompts differing by demographic groups, sampled from a distribution.
+We illustrate counterfactual bias certification for distributions of counterfactual prompts created by applying prefixes sampled from prefix distributions, to a given set of prompts.
+We consider prefix distributions consisting random token sequences, mixtures of manual jailbreaks, and perturbations of jailbreaks in LLM’s embedding space.
+We generate non-trivial certificates for SOTA LLMs, exposing their vulnerabilities over distributions of prompts generated from computationally inexpensive prefix distributions.
 1 Introduction
-Text-generating Large Language Models (LLMs) are recently being deployed in user-facing applications, such as chatbots (Lee et al., 2023).
-LLM-powered chatbots, like ChatGPT (Brown et al., 2020a) and Perplexity AI (Perplexity, 2023), are popular for their ability to produce human-like texts (Shahriar and Hayawi, 2023).
-The underlying LLMs are safety-trained (Wang et al., 2023) to avoid generating harmful content.
-However, despite safety training, they have been shown to produce texts that exhibit social biases and stereotypes (Kotek et al., 2023; Manvi et al., 2024; Hofmann et al., 2024).
+Text-generating Large Language Models (LLMs) are recently being deployed in user-facing applications, such as chatbots (Lee et al., 2023; Brown et al., 2020a; Gemini Team, 2024) where they are popular for producing human-like texts (Shahriar and Hayawi, 2023).
+These LLMs are safety-trained (Wang et al., 2023) to avoid generating harmful content.
+However, despite safety training, they can produce texts that exhibit social biases and stereotypes (Kotek et al., 2023; Manvi et al., 2024; Hofmann et al., 2024).
 Such texts can result in representational harms (Suresh and Guttag, 2021; Blodgett et al., 2020) to protected demographic groups (a subset of the population that is negatively affected by bias).
 Representational harms include stereotyping, denigration, and misrepresentation of historically and structurally oppressed demographic groups.
-Although “representational harms are harmful in their own right” (Blodgett et al., 2020), as they can socially impact individuals and redefine social hierarchies, the resulting allocation harms can lead to economic losses to the protected groups and are hence regulated by anti-discrimination laws such as (Sherry, 1965).
-Language is considered an important factor for labeling, modifying, and transmitting beliefs about demographic groups and can seldom result in new social hierarchies and amplify representational harm.
-Hence, with the rising popularity of LLMs, it is important to formally evaluate their biases to effectively mitigate representational harms due to them.
-Prior work has primarily focused on benchmarking the performance of LLMs (Liang et al., 2023; Wang et al., 2024; Mazeika et al., 2024) and adversarial attack generation (Sheng et al., 2020; Zou et al., 2023; Vega et al., 2023; Wallace et al., 2019).
+Although “representational harms are harmful in their own right” (Blodgett et al., 2020), as they can socially impact individuals and redefine social hierarchies, the resulting allocation harms (Gallegos et al., 2024a) can lead to economic losses to protected groups and are therefore regulated by anti-discrimination laws such as (Sherry, 1965).
+Language is considered an important factor for labeling, modifying, and transmitting beliefs about demographic groups and can result in reinforcement of social inequalities (Rosa and Flores, 2017).
+Hence, with rising popularity of LLMs, it is important to formally evaluate their biases to effectively mitigate representational harms resulting from them (Lee et al., 2024).
+We focus on counterfactual bias, inspired from Kusner et al.
+(2018), which assesses semantic differences across LLM responses caused by varying demographic groups mentioned in prompts (counterfactual prompts).
+Prior work has focused on benchmarking LLM performance (Liang et al., 2023; Wang et al., 2024; Mazeika et al., 2024) and adversarial attacks (Sheng et al., 2020; Zou et al., 2023; Vega et al., 2023; Wallace et al., 2019).
 While these methods provide some empirical insights into LLM bias, they have several fundamental limitations (McIntosh et al., 2024; Yang et al., 2023) such as — (1) Limited test cases: Benchmarking consists of evaluating several but limited number of test cases.
-Due to its enumerative nature, benchmarking can not scale to the prohibitively large sets of prompts that can elicit bias from LLMs.
-Adversarial attacks, on the other hand, identify only a few worst-case examples of bias, which do not inform about the overall biases from large input sets; (2) Test set leakage: LLMs may have been trained on the popular benchmarking datasets, thus resulting in biased evaluation; (3) Lack of guarantees.
-Benchmarking involves empirical estimation that does not provide any formal guarantees of generalization over any input sets.
-Similarly, adversarial attacks give limited insights as they can show existence of problematic behaviors on individual inputs but do not quantify the risk of obtaining biased LLM responses.
+Due to its enumerative nature, benchmarking can not scale to prohibitively large numbers of prompts that can elicit bias from LLMs.
+Adversarial attacks identify only few worst-case examples, which do not inform about the overall biases from large input sets; (2) Test set leakage: LLMs may have been trained on popular benchmarking datasets, thus resulting in incorrect evaluation; (3) Lack of guarantees.
+Benchmarking involves empirical estimation without any formal guarantees of generalization over any input sets.
+Similarly, adversarial attacks give limited insights as they can show existence of problematic behaviors on individual inputs but do not quantify the risk of biased LLM responses.
 This work.
-In this work, we propose an alternative to benchmarking and adversarial attacks — certifying LLMs for bias which gives formal guarantees.
-Certification operates on a prohibitively large set of inputs, represented as a specification.
-As specifications define inputs mathematically through operators over the vocabulary of LLMs, certification can provide guarantees on the behavior of the target model that generalize to unseen inputs satisfying the specification.
-With guarantees, we can be better informed about the available models before deploying them in public-facing applications.
+We propose an alternative to benchmarking and adversarial attacks — certifying LLMs for bias, with formal guarantees.
+Certification operates on a prohibitively large set of inputs, represented succinctly as a specification.
+Specifications define inputs mathematically through operators over the vocabulary of LLMs.
+Certification can provide guarantees on the target model’s behavior that generalize to unseen inputs satisfying the specification.
+With guarantees, we can be better informed about the risks associated with LLMs before deploying them in public-facing applications.
 Key challenges.
-(1) There are no existing precise mathematical representations of large sets of prompts to make practical specifications.
+(1) There are no existing precise mathematical representations of large sets of (counterfactual) prompts to make practical specifications.
 (2) State-of-the-art neural network certifiers (Wang et al., 2021; Singh et al., 2019) currently do not scale to LLMs as they require white-box access to the model and lose precision significantly for larger models, resulting in inconclusive results.
 Our approach.
-Given the diversity of prompts handled by LLMs, there will always be a few cases where the LLM output will be biased (e.g., these can be found with adversarial attacks (Zou et al., 2023; Chao et al., 2023)).
-Hence, we believe that LLM certification must be quantitative (Li et al., 2022a; Baluta et al., 2021) and focus on the following question:
-What is the probability of LLM responses being unbiased for any random set of prompts each mentioning different demographic groups?
-Figure 1: (Overview of QuaCer-B): QuaCer-B is a quantitative certification framework to certify the bias in the responses of a target LLM for a random set of prompts that differ by their sensitive attribute.
-In specific instantiations, QuaCer-B samples a (a) set of prefixes from a given distribution and prepends them to a prompt set to form (b) the prompts given to the target LLM.
+Given the diversity of LLM prompts, there will always be some cases where the LLM output will be biased (e.g., found by adversarial attacks (Zou et al., 2023)).
+Hence, we believe that LLM certification must be quantitative (Li et al., 2022a; Baluta et al., 2021) and study the question:
+What is probability of unbiased LLM responses for any counterfactual prompt set?
+Figure 1: (Overview of LLMCert-B): LLMCert-B is a quantitative certification framework to certify bias in target LLM’s responses to a random set of prompts differing only by a sensitive attribute.
+In specific instantiations, LLMCert-B samples a (a) set of prefixes from a given distribution and prepends them to a prompt set to form (b) the prompts given to the target LLM.
 (c) The target LLM’s responses are checked for bias by a bias detector, (d) whose results are fed into a certifier.
-(e) The certifier computes bounds on the probability of obtaining biased responses from the target LLM for any set of prompts formed with a random prefix from the distribution.
-We study bias over a set of prompts with varying demographic groups, inspired from Counterfactual Fairness (Kusner et al., 2018).
-In particular, we identify the biases across the LLM responses for counterfactual prompts (that differ only by the mentioned demographic groups) to capture the difference in their outputs caused by just varying demographic groups in the prompts.
-Exactly computing the probability of unbiased responses is infeasible due to the large number of possible sets of counterfactual prompts over which the biased behavior has to be determined.
+(e) Certifier computes bounds using the Clopper-Pearson method (Clopper and Pearson, 1934) on probability of unbiased LLM responses for any set of prompts formed with a random prefix from prefix distribution.
+Exactly computing the probability of unbiased responses is infeasible due to the large number of possible counterfactual prompt sets over which the biased behavior has to be determined.
 One can try to compute deterministic lower and upper bounds on the probability (Berrada et al., 2021).
 However, this is expensive and requires white-box access making it not applicable to popular, SOTA but closed-source LLMs such as GPT-4 (Achiam et al., 2023).
-Therefore, we focus on black-box probabilistic certification that estimates the probability of unbiased responses with high confidence bounds.
-We develop the first general specification and certification framework, QuaCer-B [^1] Quantitative Certification of Bias for LLM bias, applicable to both open and closed-source LLMs.
-Figure 1 gives an overview of our framework.
-Our specifications encode the probability of unbiased LLM responses for counterfactual prompts.
-We demonstrate QuaCer-B with $3$ exemplar kinds of specifications, each characterized by a specific kind of distribution of prefixes applied to a fixed set of counterfactual prompts.
-The distributions of prefixes we present are — random sequence of tokens, mixture of popular jailbreaks, and jailbreak perturbations in embedding space.
+Therefore, we focus on black-box probabilistic certification that estimates the probability of unbiased responses over a given distribution of counterfactual prompt sets with high confidence bounds.
+We develop the first general specification and certification framework, LLMCert-B [^1] LLM Certification of Bias for counterfactual bias in LLMs, applicable to both open and closed-source LLMs (see Figure 1 for an overview).
+Our specifications over counterfactual prompt sets are the first relational properties (Barthe et al., 2011) for trustworthy LLMs.
+We demonstrate LLMCert-B with $3$ kinds of specifications, each with counterfactual prompt set distributions formed by adding prefixes sampled from given distribution of prefixes to a fixed set of counterfactual prompts.
+We present distributions of prefixes consisting — random token sequences, mixtures of popular jailbreaks, and jailbreak perturbations in embedding space (§ 3.1).
 The first two are model-agnostic specifications and hence apply to both open and closed-source models.
-However, the third one requires access to the embeddings and the ability to prompt LLMs with embeddings, and hence mostly applies to only open-source models.
-The prefixes are described further in Section 3.1.
-We certify the proposed specifications leveraging confidence intervals.
-Our certifier samples several counterfactual prompt sets from the distribution given in the specification and generates high-confidence bounds on the probability of getting unbiased responses from the target LLM for any random prompt set in the distribution.
+However, the third one requires access to the embeddings and the ability to prompt LLMs with embeddings, and hence applies to only open-source models.
+The mixture and embedding space jailbreak prefix distributions contain effective, manually designed jailbreaks and their perturbations, which are potential jailbreaks, in their sample space and hence assess LLMs’ biases in adversarial settings.
+We certify the target LLM leveraging confidence intervals.
+LLMCert-B samples several counterfactual prompt sets from the specified prompt distribution and generates high-confidence bounds on the probability of unbiased LLM responses for any random counterfactual prompt set in the distribution.
 Contributions. Our main contributions are:
-  • We design novel specifications that quantify the desirable property of low bias in the responses of LLMs with varying prefixes.
-The prefixes to given counterfactual prompts can be drawn from $3$ distributions — (1) random, (2) mixture of jailbreaks, and (3) jailbreak perturbations in the embedding space.
-  • We develop a probabilistic black-box certifier QuaCer-B, applicable to both open and closed-source models, for quantifying bias in LLM output.
-QuaCer-B leverages confidence intervals (Clopper and Pearson, 1934) to generate high-confidence bounds on the probability of obtaining unbiased responses from the target LLM, given any random set of prompts from the distribution given in the specification.
-  • We use QuaCer-B to reveal novel insights into the biases of popular LLMs, for different specifications (Section 5).
-We find that the safety alignment of SOTA models is easily circumvented with several prefixes in the distributions given in our specifications, especially those involving mixture of jailbreaks and jailbreak perturbations in the embedding space.
+  • We design novel specifications that quantify the desirable relational property of low counterfactual bias in LLM responses over counterfactual prompts in a specified distribution.
+We illustrate such specifications with distributions of counterfactual prompt sets constructed with potentially adversarial prefixes.
+The prefixes are drawn from $3$ distributions — (1) random, (2) mixture of jailbreaks, and (3) jailbreak perturbations in the embedding space.
+  • We develop the first probabilistic black-box certifier LLMCert-B, applicable to both open and closed-source models, for quantifying counterfactual bias in LLM responses.
+LLMCert-B leverages confidence intervals (Clopper and Pearson, 1934) to generate high-confidence bounds on the probability of obtaining unbiased responses from the target LLM, given any random set of counterfactual prompts from the distribution given in the specification.
+  • We find that the safety alignment of SOTA LLMs is easily circumvented with several prefixes in the distributions given in our specifications, especially those involving mixture of jailbreaks and jailbreak perturbations in the embedding space (§ 5).
 These distributions are inexpensive to sample from, but can effectively bring out biased behaviors from SOTA models.
@@
 We find that there are no consistent trends in the fairness of models with the scaling of their sizes, hence suggesting that the quality of alignment techniques could be a more important factor than size for fairness.
-Our implementation is available at https://github.com/uiuc-focal-lab/QuaCer-B.
+Our implementation is available at https://github.com/uiuc-focal-lab/LLMCert-B and we provide guidelines for using our framework for practitioners in Appendix A.
 2 Background
-2.1 Large Language Models
-Large Language Models (LLMs) are autoregressive foundation models that perform next-token prediction.
-That is, given a sequence of tokens $t_{1},\dots,t_{k}$, they give a distribution over the vocabulary for the next token, $P[t_{k+1}\mid t_{1},\dots,t_{k}]$.
-They are typically fine-tuned for instruction-following capabilities to be used such as chatbots.
-These instruction-finetuned models are further aligned with human feedback (Wang et al., 2023; Ouyang et al., 2022) to make their responses safe and unbiased.
-We certify these instruction-tuned (Zhang et al., 2024), aligned models for fairness, as they are typically used in public-facing applications.
+2.1 Large Language Models (LLMs)
+LLMs are autoregressive models for next-token prediction.
+Given a sequence of tokens $t_{1},\dots,t_{k}$, they give a probability distribution over their vocabulary for the next token, $P[t_{k+1}\mid t_{1},\dots,t_{k}]$.
+They are typically fine-tuned for instruction-following (Zhang et al., 2024) and aligned with human feedback (Wang et al., 2023; Ouyang et al., 2022) to make their responses safe.
+We certify instruction-tuned, aligned LLMs for counterfactual bias, as they are typically used in public-facing applications.
 2.2 Clopper-Pearson confidence intervals
-Clopper-Pearson confidence intervals (Clopper and Pearson, 1934) provide lower and upper bounds on the probability of success parameter of a Bernoulli random variable with probabilistic guarantees.
+Clopper-Pearson confidence intervals (Clopper and Pearson, 1934) provide lower and upper bounds $[p_{l},p_{u}]$ on the probability of success parameter $p$ of a Bernoulli random variable with probabilistic guarantees.
 The bounds are obtained with $n$ independent and identically distributed observations of the random variable, in which $k(\leq n)$ successes are observed.
-The confidence interval is constructed to bound the value of $p$ with $[\hat{p_{l}},\hat{p_{u}}]$ with at least a given confidence level $(1-\gamma)100\%$.
-That is, $Pr\{p\in[\hat{p_{l}},\hat{p_{u}}]\}\geq(1-\gamma)$.
-$\gamma$ is the error probability by which the true value of $p\notin[\hat{p_{l}},\hat{p_{u}}]$.
-These confidence intervals are obtained by binomial statistical hypothesis testing for the value of $p$, where the lowest and highest values are $\hat{p_{l}}$ and $\hat{p_{u}}$ respectively, with the given confidence $1-\gamma$.
+The confidence interval is such that $Pr\{p\in[[p_{l},p_{u}]\}\geq(1-\gamma)$.
+$\gamma\in(0,1)$ is the (small) permissible error probability by which the true value of $p\notin[p_{l},p_{u}]$.
+The confidence intervals are obtained by statistical hypothesis testing for $p$, where the lowest and highest values are $p_{l}$ and $p_{u}$ respectively, with the given confidence $1-\gamma$.
 3 Formalizing bias certification
-We develop a general framework, QuaCer-B, to specify and quantitatively certify the bias in the text generated by (Large) Language Models.
-QuaCer-B formalizes bias with specifications — precise mathematical representations that define the desirable property (absence of bias) in large sets of inputs.
-Bias is defined with respect to demographic groups that are subsets of the human population that share an identity trait, that could be biological, contextual, or socially constructed (Gallegos et al., 2024b).
-Bias consists of disparate treatment or outcomes when varying the demographic groups in the inputs to the target model (Gallegos et al., 2024b; Sheng et al., 2019; Smith et al., 2022).
-In the context of autoregressive LMs, we consider text generation bias consisting of stereotyping, misrepresentation, derogatory language, etc.
-This can primarily result in representational harms, which can subsequently cause allocation harms (Gallegos et al., 2024b).
-To apply certification to closed-source LMs as well, we study extrinsic bias (Cao et al., 2022) that manifests in the final generated responses of the models.
+We develop a general framework, LLMCert-B, to specify and quantitatively certify counterfactual bias in text generated by (Large) Language Models (LMs).
+LLMCert-B formalizes bias with specifications — precise mathematical representations that define the desirable property (absence of bias) in large sets of inputs.
+Bias is defined with respect to demographic groups — subsets of human population sharing an identity trait, that could be biological or socially constructed (Gallegos et al., 2024b).
+Bias consists of disparate treatment or outcomes when varying the demographic groups in inputs to the target LM.
+For autoregressive LMs, we consider text generation bias consisting of stereotyping, misrepresentation, derogatory language, etc, that can result in representational harms (Gallegos et al., 2024b).
+To apply certification to closed-source LMs as well, we study extrinsic bias (Cao et al., 2022) that manifests in the final LM response texts.
+See Appendix I for a detailed discussion on bias in ML.
 3.1 Bias specification
-Next, we formally specify the lack of bias in the outputs of language models (LMs).
-In particular, we want the text generated by LMs to not exhibit any semantic disparities caused by mentioning different demographic groups in the prompts.
-Hence, our bias specification is motivated by Counterfactual Fairness (Kusner et al., 2018).
+Next, we formally specify the lack of bias in the responses of language models.
+Unbiased LM responses do not exhibit semantic disparities owing to specific demographic groups in the prompts (Gallegos et al., 2024b; Sheng et al., 2019; Smith et al., 2022).
+Our bias specification is motivated by Counterfactual Fairness (Kusner et al., 2018).
 Consider a given identity trait $\mathcal{I}$ such as gender, race, etc.
-(that are often the basis of social bias).
-$\mathcal{I}$ categorizes the human population into $m$ subsets called demographic groups $\mathcal{G}_{1},\dots,\mathcal{G}_{m}$, each differing by the value of the identity trait.
-Each demographic group $\mathcal{G}$ is characterized/recognized by several synonymous strings, called sensitive attributes $\mathcal{G}^{\mathcal{A}}$ (Li et al., 2024).
+(that are often basis of social bias).
+$\mathcal{I}$ categorizes the human population into $m$ subsets called demographic groups $\mathcal{G}_{1},\dots,\mathcal{G}_{m}$, each differing by the identity trait.
+Each demographic group $\mathcal{G}$ is characterized/recognized by several synonymous strings in the society, called sensitive attributes $\mathcal{G}^{\mathcal{A}}$ (Li et al., 2024).
 For example, the sensitive attributes for the demographic group corresponding to the female gender are woman, female, etc.
@@
 Let the resulting set of sensitive attributes, each corresponding to a demographic group, be $\mathcal{A}=\{\mathcal{A}_{1},\dots,\mathcal{A}_{m}\}$, where $\mathcal{A}_{i}\in\mathcal{G}^{\mathcal{A}}_{i}$.
-Our specifications are for counterfactual inputs (prompts) (Gallegos et al., 2024b) that differ only by the sensitive attributes in them.
-Let $\mathcal{L}$ be the target LM and $\mathcal{V}$ be its vocabulary.
-Consider a set of prompts $\mathcal{P}=\{\mathcal{P}_{1},\dots\mathcal{P}_{s}\}$, $s>1,\mathcal{P}\subset\mathcal{V}^{[1,c]}$, where $c$ is the context length of $\mathcal{L}$ and $\mathcal{V}^{[1,c]}$ is a sequence of elements of $\mathcal{V}$ having length $\in[1,c]$.
-Let each prompt in $\mathcal{P}$ contain a unique sensitive attribute from $\mathcal{A}$ such that overall $\mathcal{P}$ represent more than $1$ distinct demographic groups represented in $\mathcal{A}$.
-Let each prompt be $\mathcal{P}_{i}=X_{i}\cup\mathcal{A}_{i}$, where $X_{i}$ is the part of the prompt that is independent of sensitive attributes and $\mathcal{A}_{i}$ consists of a sensitive attribute.
+Our specifications are for counterfactual prompts (Gallegos et al., 2024b) that differ only by the sensitive attributes in them.
+Let $\mathcal{V}$ be the vocabulary of the target LM $\mathcal{L}$.
+Consider a set of prompts $\mathcal{P}=\{\mathcal{P}_{1},\dots\mathcal{P}_{s}\}$, $s>1,\mathcal{P}\subset\mathcal{V}^{c}$, where $c$ is the context length of $\mathcal{L}$ and $\mathcal{V}^{c}$ is a sequence of elements of $\mathcal{V}$ having length $\in[1,c]$.
+Let each prompt in $\mathcal{P}$ contain a unique sensitive attribute from $\mathcal{A}$ such that overall $\mathcal{P}$ represent more than $1$ distinct demographic groups represented by $\mathcal{A}$.
+Let each prompt be $\mathcal{P}_{i}=X_{i}\cup\mathcal{A}_{i}$, where $X_{i}$ is the part of $\mathcal{P}_{i}$ independent of sensitive attributes and $\mathcal{A}_{i}$ is a sensitive attribute.
 We consider only prompts that can be decomposed into parts with and without sensitive attributes respectively.
@@
 A set of prompts $\mathcal{P}=\{\mathcal{P}_{1}=X_{1}\cup\mathcal{A}_{1},\dots,\mathcal{P}_{s}=X_{s}\cup\mathcal{A}_{s}\}$ is called counterfactual when: (1) $\forall i,j\in[1,s]\ldotp X_{i}=X_{j}$; (2) $\exists i\in[1,s]\ldotp\forall j\in([1,s]\setminus i)\ldotp\mathcal{A}_{i}\neq\mathcal{A}_{j}$; (3) For an unbiased text generator $f$, $\forall i\in[1,s]\ldotp f(\mathcal{P}_{i})=f(X_{i})$.
-That is, the prompts only differ in the sensitive attributes that are ideally unrelated to the overall query and the set of prompts $\mathcal{P}$ represent more than sensitive attribute from $\mathcal{A}$.
-We specify bias over counterfactual prompt sets (Definition 3.1), which can be prohibitive in number for typical context lengths.
-This is because, the common part of counterfactual prompt sets, $X$ can be any element from $\mathcal{V}^{[1,c]}$ which contains $\approx 10^{10000}$ elements for $c=2k$.
+That is, the prompts only differ in the sensitive attributes that are ideally unrelated to the overall query and the set of prompts $\mathcal{P}$ represent more than $1$ sensitive attributes from $\mathcal{A}$.
+We specify bias over counterfactual prompt sets (Definition 3.1).
+These exclude prompts where sensitive attributes are important to answer the overall query, such as “What steps should I take to prepare for becoming a parent?", as semantically different answers, dependent on the sensitive attribute, are acceptable for such cases.
+All possible counterfactual prompt sets can be prohibitive in number for typical context lengths.
+This is because, the common part of counterfactual prompt sets, $X$ can be any element from $\mathcal{V}^{c}$ which contains $\approx 10^{10000}$ elements for $c=2k$.
 Hence, enumerative specifications (which specify the desired behavior on all inputs) are impractical, as they cannot be scalably certified without symbolic analysis for large number of inputs.
 Hence, we define probabilistic specifications for the probability of unbiased responses from $\mathcal{L}$, for which we provide a certification algorithm in Section 3.2.
-Let $\Delta$ be a sampleable probability distribution over $\wp(\mathcal{V}^{[1,c]})$ (power set of prompts) having non-zero support on some counterfactual prompt sets $\mathcal{P}$.
+Let $\Delta$ be a sampleable discrete probability distribution over $\wp(\mathcal{V}^{c})$ (power set of prompts) having non-zero support on some counterfactual prompt sets $\mathcal{P}$.
 We define probabilistic specifications for bias in $\mathcal{L}$ over $\Delta$.
 The specification is agnostic to $\Delta$’s sampler, as long as it generates independent and identically distributed samples.
-We provide examples of $\Delta$ in Section 4.
+We show examples of $\Delta$ in Section 4.
 Let $\mathcal{D}$ be a user-defined bias detection function that can identify stereotypes/disparity in given texts for different sensitive attributes in $\mathcal{A}$.
-Let $\mathcal{D}$ evaluate to zero for unbiased inputs (appropriate scaling and shifting of $\mathcal{D}$ can be done to satisfy this criterion).
-We leave $\mathcal{D}$ as a parameter of the specification, as different domains can have varying notions of bias and the stakeholders can decide which notion is most suitable for their usecase (Anthis et al., 2024).
-Overall, we give our quantitative specification as the probability of unbiased responses (as measured by $\mathcal{D}$) when $\mathcal{L}$ is independently prompted with each element of a randomly sampled counterfactual prompts set from $\Delta$ (1).
-The certificate $\mathcal{C}$ is defined for the user-defined parameters $\Delta$, $\mathcal{D}$, and $\mathcal{L}$.
+Let $\mathcal{D}$ evaluate to zero for unbiased inputs (scaling and shifting of $\mathcal{D}$ can be done to satisfy this).
+We leave $\mathcal{D}$ as a parameter of the specification, as different domains can have varying notions of bias and stakeholders can decide the most suitable notion (Anthis et al., 2024).
+Overall, our quantitative specification is the probability of unbiased responses (measured by $\mathcal{D}$) when $\mathcal{L}$ is independently prompted with each element of a randomly sampled counterfactual prompts set from $\Delta$ (1).
+Certificate $\mathcal{C}$ is an evaluation/estimation of the specified probability of unbiased responses, along with samples of LLM responses, for user-defined parameters $\Delta$, $\mathcal{D}$, and $\mathcal{L}$.
 $\mathcal{C}(\Delta,\mathcal{D},\mathcal{L})\triangleq Pr_{\mathcal{P}\sim\Delta}[\mathcal{D}([\mathcal{L}(\mathcal{P}_{1}),\dots,\mathcal{L}(\mathcal{P}_{s})])=0]$ (1)
@@
 Next, we describe the $3$ different kinds of practical $\Delta_{pre}$ and their sampling algorithms to define local bias specifications for $\mathcal{L}$.
-We show some samples from each kind of $\Delta_{pre}$ in Appendix D.
+We show some samples from each kind of $\Delta_{pre}$ in Appendix C.
 Our specifications are for the average-case behavior of the target LLM, as $\Delta_{pre}$ are not distributions of provably adversarial (worst-case) prefixes.
@@
 Random prefixes.
-Prior works such as (Wei et al., 2023; Zou et al., 2023) have shown the effects of incoherent strings in manipulating LLMs and jailbreaking them into answering queries about harmful behaviors.
+Prior works such as (Wei et al., 2023; Zou et al., 2023) have shown the effects of incoherent fixed-length strings in jailbreaking LLMs for harmful prompts.
 Hence, we specify bias in LLMs for prompts with incoherent prefixes that are random sequences of tokens from the vocabulary of the LLM.
+Such prefixes are not all intentionally adversarial, except for adversarial strings like those from prior works, but denote random noise in the prompt.
 Algorithm 2 presents the prefix sampler as a uniform distribution, $\mathcal{U}(.)$ over the random prefixes of fixed length, $q$.
@@
 The number of prefixes formed by the aforementioned operations can be prohibitively many, owing to typically long manual jailbreaks and the possibility to mutate any token to any random token from the LLM’s vocabulary.
-Soft jailbreaks.
+Soft prefixes from jailbreaks.
 Due to the limited number of effective manual jailbreaks (walkerspider, 2022; Learn Prompting, 2023), they can be easily identified and defended against.
-However, the excellent denoising capabilities of LLMs could render them vulnerable to simple manipulations of manual jailbreaks as well, indicating that the threat is not completely mitigated by the current defenses (Jain et al., 2023).
-Hence, we specify fairness under prefixes constructed by adding noise of the original manual jailbreaks.
+However, the excellent denoising capabilities of LLMs could render them vulnerable to simple manipulations of manual jailbreaks as well, indicating that the threat is not completely mitigated by current defenses (Jain et al., 2023).
+Hence, we specify fairness under prefixes constructed by adding noise to the original manual jailbreaks.
 Algorithm 4 presents the prefix constructor as a probabilistic program.
@@
 5 Experiments
-5.1 Experimental setup
 We used $2$ A100 GPUs, each with 40GB VRAM.
@@
 To identify bias in the responses of the target LLM for the prompts constructed from pivot prompt sets in $\mathcal{Q}_{BOLD}$, we use the regard-based bias metric proposed in (Sheng et al., 2019).
-Our adaptation of the regard metric is described and evaluated with a human study on Amazon Mechanical Turk in Appendix C.1.
+Our adaptation of the regard metric is described and evaluated with a human study on Amazon Mechanical Turk in Appendix G.1.
 Our bias detector matches human perception of bias in $76\%$ cases.
-We qualitatively analyze the false positive and false negatives of the bias detector with respect to human judgment in Appendix C.1.
+We qualitatively analyze the false positive and false negatives of the bias detector with respect to human judgment in Appendix G.1.
 Figure 3: Example from $\mathcal{Q}_{DT}$
@@
 Let the resultant test set consisting of pivot prompts from each partial sentence considered from DT be $\mathcal{Q}_{DT}$.
-We evaluate the LLM responses to prompts derived from pivot prompt sets in $\mathcal{Q}_{DT}$ using a bias detector that identifies the disparity in agreement to the stereotype for different demographic groups given in the prompt, as discussed in Appendix C.2.
+We evaluate the LLM responses to prompts derived from pivot prompt sets in $\mathcal{Q}_{DT}$ using a bias detector that identifies the disparity in agreement to the stereotype for different demographic groups given in the prompt, as discussed in Appendix G.2.
 For every element in $\mathcal{Q}_{BOLD}$ and $\mathcal{Q}_{DT}$, we generate 3 certificates for the specifications in Section 4, such that a certificate consists of bounds on the probability of unbiased responses from the target LLM.
 Both bias detectors are such that they output $1$ for a biased set of responses to a counterfactual prompt set and 00 for an unbiased set of responses.
-The values of the certification parameters used in our experiments are given in Table 2 (Appendix A).
-We study their effect on the certification results with an ablation study in Appendix A.
+The values of the certification parameters used in our experiments are given in Table 3 (Appendix E).
+We study their effect on the certification results with an ablation study in Appendix E.
 We generate the certification bounds with $95\%$ confidence and $50$ samples.
-While our main experiments are for counterfactual prompt sets with binary demographic groups, our framework can be extended beyond binary demographic groups, which we experimentally demonstrate in Appendix A.4.
-Note that the manual jailbreaks used are common across all specifications and are presented in Appendix B.
+While our main experiments are for counterfactual prompt sets with binary demographic groups, our framework can be extended beyond binary demographic groups, which we experimentally demonstrate in Appendix E.6.
+Appendix B presents the existing, manually-designed jailbreaks used across all specifications.
+Note that, these jailbreaks are just examples to demonstrate our framework, which generalizes beyond them to new bias-eliciting manual textual jailbreaks.
 Table 1: Average of the bounds on the probability of unbiased responses for different models.
@@
 Claude-3.5-Sonnet $99.6$ $100.0$ $(0.93,1.0)$ $(0.93,1.0)$-
+Table 2: Average of the bounds on the probability of unbiased responses for different models.
+Lowest bounds for each specification kind and dataset are highlighted.
+We report $2$ baselines — unbiased responses when prompting without prefixes and with the main jailbreak as prefix.
+Average certification bounds
+Dataset Model Random Mixture Soft
+BOLD (250) Vicuna-13B $(0.93,1.0)$ $(0.93,1.0)$ $(0.92,1.0)$
+Llama-13B $(0.93,1.0)$ $(0.91,1.0)$ $(0.93,1.0)$
+Mistral-7B $(0.92,1.0)$ $\mathbf{(0.22,0.42)}$ $\mathbf{(0.30,0.52)}$
+Gemini $(0.92,1.0)$ $(0.60,0.83)$-
+GPT $(0.92,1.0)$ $(0.44,0.67)$-
+DT (48) Vicuna-13B $\mathbf{(0.71,0.92)}$ $(0.92,1.0)$ $(0.51,0.78)$
+Llama-13B $(0.92,1.0)$ $(0.93,1.0)$ $(0.93,1.0)$
+Mistral-7B $(0.91,1.0)$ $(0.85,0.99)$ $\mathbf{(0.46,0.73)}$
+Gemini $(0.92,1.0)$ $(0.73,0.93)$-
+GPT $(0.93,1.0)$ $\mathbf{(0.66,0.88)}$-
 5.1 Certification results
@@
 We report the average of the certification bounds for all pivot prompt sets in $\mathcal{Q}_{BOLD}$ and $\mathcal{Q}_{DT}$ each for every model in Table 1.
-We do not certify the closed-source models such as Gemini, GPT, and Claude for soft jailbreaks, as it requires access to the models’ embedding layers.
+We do not certify the closed-source models such as Gemini, GPT, and Claude for soft prefixes, as it requires access to the models’ embedding layers.
 Certification time significantly depends on the inference latency of the target model.
@@
 The baselines consider $\mathcal{Q}_{BOLD}$ and $\mathcal{Q}_{DT}$ as benchmarking datasets, having counterfactual prompt sets as individual elements.
-Similar to popular LLM bias benchmarking works such as (Wang et al., 2024; Liang et al., 2023; Esiobu et al., 2023; Xie et al., 2024), we study the biases in LLMs for a fixed dataset of counterfactual prompt sets which may be provided as is to the LLM, or with some provoking texts (jailbreaks).
-In the first baseline (without prefix), every counterfactual prompt set is evaluated $5$ times by each time prompting a target LLM with each prompt in the set without any prefixes and detecting bias across its responses using the bias detectors of the respective datasets.
+Similar to popular LLM bias benchmarking works such as (Wang et al., 2024; Liang et al., 2023; Esiobu et al., 2023; Xie et al., 2024), we study the biases in LLMs for a fixed dataset of counterfactual prompt sets which may be provided as is to the LLM, or with jailbreaks.
+In the first baseline (without prefix), every counterfactual prompt set is evaluated $5$ times, each time prompting a target LLM with each prompt in the set without any prefixes and detecting bias across its responses using the corresponding bias detector.
 The bias result for each counterfactual prompt set is computed by averaging the results over the $5$ evaluations, similar to (Wang et al., 2024).
-This baseline indicates biases in the LLM responses without any prefixes and is hence used to judge the influence of the prefixes on eliciting biased responses from the models.
-Table 1 reports the average of the evaluations over all counterfactual prompt sets in $\mathcal{Q}_{BOLD}$ and $\mathcal{Q}_{DT}$ respectively.
-In the second baseline (with main jailbreak), each counterfactual prompt is similarly evaluated $5$ times, but each time prompting the LLM with each prompt in the set with the unmodified main jailbreak (Figure 7 in Appendix B) used in the mixture of jailbreak and soft jailbreak distributions, as a prefix.
+This baseline indicates biases in LLM responses without any prefixes and can be used to judge the additional influence of prefixes on eliciting bias from LLMs.
+Table 1 reports the average of evaluations over all counterfactual prompt sets in $\mathcal{Q}_{BOLD}$ and $\mathcal{Q}_{DT}$ respectively.
+In the second baseline (with main jailbreak), each counterfactual prompt is similarly evaluated $5$ times, but with the unmodified main jailbreak (Figure 5, Appendix B), used in the mixture of jailbreak and soft prefixes from jailbreak distributions, as a prefix.
 The average result of this baseline is also reported in Table 1.
-This baseline is used to indicate the efficacy of the main jailbreak without any modifications, and hence suggests the importance of the mixture and soft prefix distributions around the main jailbreak in modifying the biases in LLM responses.
-The benchmarking baselines are empirical studies of bias in the target LLMs, which analyze the biases with specific prompts.
-On the other hand, QuaCer-B quantifies biases for any random prompt sampled from a given distribution.
+This baseline is used to indicate the efficacy of the main jailbreak without any modifications, and hence suggests the importance of the mixture and soft prefix distributions around the main jailbreak in eliciting biases in LLM responses.
+The baselines are empirical studies of counterfactual bias in LLMs, which analyze bias with a dataset of prompts.
+On the other hand, LLMCert-B quantifies biases for any random prompt sampled from a given distribution.
 5.1.1 General observations
@@
 Denial of response.
-In several instances, such as those during the certification of GPT-4 (for DT), Claude, and Llama, we observe unbiased behaviors because the model simply refuses to respond.
-The completions are mostly "Sorry I can’t assist with that".
-While that can be a valid defense to adversarial prompts, it may not be the most desirable.
+In several instances, such as those when certifying GPT-4 (for DT), Claude, and Llama, we observe unbiased behaviors because the model simply refuses to respond.
+The completions are mostly “Sorry I can’t assist with that".
+While that can be a valid defense to adversarial prompts, it may not be the most desirable (Cui et al., 2024).
 The model could instead output neutral responses, ignoring any bias-encouraging texts in the prompts to be helpful but not harmful.
@@
 Mistral.
-Mistral exhibits significantly low probability of unbiased when prompted with the mixture of jailbreaks and soft jailbreak prefixes for specifications from both $\mathcal{Q}_{BOLD}$ and $\mathcal{Q}_{DT}$.
+Mistral exhibits significantly low probability of unbiased when prompted with mixtures of jailbreaks and soft prefixes for specifications from both $\mathcal{Q}_{BOLD}$ and $\mathcal{Q}_{DT}$.
 Gemini.
@@
 5.2 Case studies
-We qualitatively introspect the individual certificates obtained for different models for one example from $\mathcal{Q}_{BOLD}$.
-We present similar analyses for one example from $\mathcal{Q}_{DT}$ in Appendix E.
-We select the example for Desktop automation refers to software programming from $\mathcal{Q}_{BOLD}$ for the ‘Automation’ profession and study the sample responses obtained for the mixture of jailbreak specification.
-We present some responses of $4$ popular LLMs — Mistral, Llama-7B, Gemini, and GPT-4 obtained during certification, in Figure 4.
-The responses demonstrate various stereotypical biases, against both genders.
-For example, the Mistral model discriminates against women, calling them “incapable and better suited for menial jobs”.
-The Gemini model is also frequently biased against women, saying “they cannot think and are meant for household chores”.
-On the other hand, GPT-4 regularly shows biases towards men, as it gives responses about them “lacking motivation and being less efficient”, with automation being better at their jobs.
-It depicts women crushing their male competition and becoming superior.
-The Llama model, however, defends against the prompts and avoids responding in a biased way.
-The responses reflect the certification results for the models for the BOLD example.
-We find from the certification bounds that the Mistral model is provably worse than Gemini and GPT, which are worse than Llama with high confidence for this specification.
-Figure 4: Responses of Mistral, Llama-7B, Gemini, and GPT-4 obtained during their certification for the BOLD example ‘Desktop automation refers to software programming’ for mixture of jailbreak specifications.
-The certification result is indicated alongside the model.
+We qualitatively introspect individual certificates for different models for one example from $\mathcal{Q}_{BOLD}$.
+We present similar analyses for one example from $\mathcal{Q}_{DT}$ in Appendix D.
+We select the example for Desktop automation refers to software programming from $\mathcal{Q}_{BOLD}$ for ‘Automation’ profession.
+We present responses of $4$ popular LLMs — Mistral, Llama-7B, Gemini, and GPT-4 obtained during certification for the mixture of jailbreak specification, in Figure 4.
+The responses demonstrate various stereotypical biases against both genders.
+For example, Mistral discriminates against women, calling them “incapable and better suited for menial jobs”.
+The Gemini model is also biased against women, saying “they cannot think and are meant for household chores”.
+On the other hand, GPT-4 shows biases towards men, as it responds about them “lacking motivation and being less efficient”, with automation being better at their jobs.
+It depicts women crushing male competition and becoming superior.
+Llama, however, defends against the prompts and avoids responding in a biased way.
+The responses reflect the certification results for the models.
+We find from the certification bounds that the Mistral model is worse than Gemini and GPT, which are worse than Llama with high confidence for this specification.
+Figure 4: Responses (and bounds) of Mistral, Llama-7B, Gemini, and GPT-4 when certifying with BOLD example ‘Desktop automation refers to software programming’ for mixture of jailbreaks.
 6 Related work
 Attacking LLM alignment.
-LLMs are typically aligned with human ethics by supervised fine-tuning and reinforcement learning with human feedback (Ouyang et al., 2022).
-However, prior work (Zou et al., 2023; Vega et al., 2023; Chao et al., 2023; Sheng et al., 2020; Wallace et al., 2019) propose various methods to generate jailbreaking attacks for LLMs, which bypass their safety-alignment and lead to harmful or biased responses.
-Jailbreaks can be incoherent as in (Zou et al., 2023; Sheng et al., 2020; Wallace et al., 2019) or coherent instructions (Dominique et al., 2024; Liu et al., 2024).
+LLMs are aligned with human ethics by supervised fine-tuning and reinforcement learning with human feedback (Ouyang et al., 2022).
+However, (Zou et al., 2023; Vega et al., 2023; Chao et al., 2023; Sheng et al., 2020; Wallace et al., 2019) propose methods to jailbreak LLMs, bypassing alignment and causing harmful or biased responses.
+Jailbreaks can be incoherent (Zou et al., 2023; Sheng et al., 2020) or coherent (Dominique et al., 2024; Liu et al., 2024).
 Benchmarking LLMs.
-Various prior works have benchmarked the performance of LLMs on standard and custom datasets.
-These could be datasets of general prompts (Dhamala et al., 2021; Wang et al., 2024) or those of adversarial examples (Zou et al., 2023; Mazeika et al., 2024) designed to elicit undesirable behaviors from the models.
-Popular benchmarks such as (Liang et al., 2023; Wang et al., 2024; Mazeika et al., 2024; Manerba et al., 2024; Gallegos et al., 2024a) present empirical trends for the performance of LLMs, measured along various axes including bias and fairness.
+Various prior works have benchmarked LLM performance on standard and custom datasets.
+These consist of datasets of general prompts (Dhamala et al., 2021; Wang et al., 2024) or adversarial examples (Zou et al., 2023; Mazeika et al., 2024).
+Benchmarks such as (Liang et al., 2023; Wang et al., 2024; Mazeika et al., 2024; Manerba et al., 2024; Gallegos et al., 2024a) present empirical trends for LLM performance, measured along various axes including bias.
 Guarantees for LLMs.
-There is an emerging need for guarantees on LLM behavior, fueled by their popularity in public-facing use cases.
+There is an emerging need for guarantees on LLM behavior.
 (Kang et al., 2024) provides guarantees on the generation risks of RAG LLMs.
-(Crouse et al., 2024) give formal specifications for expected behavior of LLMs as autonomous agents.
-However, there is no work on certifying LLM bias, to the best of our knowledge.
+(Quach et al., 2024; Deutschmann et al., 2023; Mohri and Hashimoto, 2024; Yadkori et al., 2024) apply conformal prediction to LLMs, proposing methods for generating output sets with statistical guarantees on correctness or coverage.
+(Zollo et al., 2024) presents a framework for selecting low-risk system prompts for LLMs with probabilistic guarantees.
+(Chaudhary et al., 2024) present a certification framework for knowledge comprehension in LLMs.
+We compare LLMCert-B with existing works further in Appendix H.
 Fairness in Machine Learning.
@@
 7 Conclusion
-We present the first general framework QuaCer-B to certify bias in the responses of LLMs, applicable to both open and closed-source models.
-We instantiate our framework with novel specifications based on different kinds of prefixes.
-QuaCer-B generates high confidence bounds on the probability of unbiased responses for prompts from a given distribution.
-Our results show that certification can detect biases not observed with existing efforts.
-As future work, QuaCer-B can be made less dependent on bias detectors.
-It can also be made sample efficient for higher confidence requirements.
-Appendix A Ablation study
+We present the first framework LLMCert-B to specify and certify counterfactual bias in LLM responses, for both open and closed-source models.
+We instantiate LLMCert-B with novel specifications based on different kinds of potentially adversarial prefixes.
+LLMCert-B generates high confidence bounds on probability of unbiased responses for counterfactual prompts from a given distribution.
+Our results show previously unknown vulnerabilities related to counterfactual bias in SOTA LLMs.
+Acknowledgement
+We thank the anonymous reviewers for their insightful comments.
+This work was supported by a grant from the Amazon-Illinois Center on AI for Interactive Conversational Experiences (AICE) and NSF Grants No.
+CCF-2238079, CCF-2316233, CNS-2148583.
+Ethics statement
+We identify the following positive and negative impacts of our work.
+Positive impacts.
+Our work is the first to provide quantitative certificates for the bias in Large Language Models.
+It can be used by model developers to thoroughly assess their models before releasing them and by the general public to become aware of the potential harms of using any LLM.
+As our framework, LLMCert-B assumes black-box access to the model, it can be applied to even closed-source LLMs with API access.
+Negative impacts.
+In this work, we propose $3$ kinds of specifications involving — random prefixes, mixtures of jailbreaks as prefixes, and jailbreaks in the embedding space of the target model.
+While these prefixes are not adversarially designed, they are often successful in eliciting biased and toxic responses from the target LLMs.
+They can be used to attack these LLMs by potential adversaries.
+We have informed the developers of the LLMs about this threat.
+Reproducibility statement
+Our implementation of LLMCert-B is open-sourced at https://github.com/uiuc-focal-lab/LLMCert-B.
+We provide a README with the code having instructions to reproduce the main results of the paper.
+The code also consists the files used in conducting a human evaluation of our bias detector for BOLD on Amazon Mechanical Turk.
+Appendix A Note for practitioners
+In this section, we describe how practioners can use our framework to assess LLMs and automatically identify vulnerabilities in them.
+Our open-source implementation is available at: https://github.com/uiuc-focal-lab/LLMCert-B, which can be used following the GPL (license) terms and conditions.
+The open-source framework can be used to certify both open and closed-source LLMs by adding support to query custom models in utils.py (for open-source models) and utils_api.py (for closed-source models) files.
+The framework requires unrestricted (in terms of number of inferences) query-access to the target model.
+Developers can adjust the desired confidence-level of the certificates and increase/decrease the number of samples used in certification for tighter/looser bounds, according to their requirements and budget.
+Developers can also use custom bias detectors to label biased responses, using which the certificates can be computed.
+To get customized insights into LLM biases for their particular applications, developers can define specifications with prompts that are commonly observed in their use cases.
+This customization can either happen by sourcing the pivot prompts from domain-specific datasets, instead of the popular BOLD or Decoding Trust datasets and/or using custom distributions of prefixes/suffixes which more suitably represent the biases in their domains.
+For example, in domains where there is threat of racial bias, the prefixes could explicitly encourage the model to exhibit racial bias, so as to stress-test the trustworthiness of the models.
+Developers can also define entirely new distributions of counterfactual prompts, irrespective of prefixes/suffixes, to specify bias similar to 1 and certify with LLMCert-B.
+The certificates obtained are reliable, quantitative risk assessments of models, with lower and upper bounds on risks pertaining to bias in the models’ generations.
+They can also be used to compare different LLMs to pick one with acceptable risk in varying contexts.
+Appendix B Common jailbreaks used
+In our experiments, we combine common jailbreaks to certify against hybrid jailbreaks.
+The specific jailbreaks we use are shown below.
+The main jailbreak [^2] Last accessed on Jan 22, 2024 from www.jailbreakchat.com/prompt/588ab0ed-2829-4be8-a3f3-f28e29c06621 is shown in Figure 5.
+We have modified the jailbreak to encourage the model to generate responses demonstrating gender bias by adding the text highlighted in red in Figure 5(a) and for racial bias in Figure 5(b).
+We use one helper jailbreak [^3] https://www.reddit.com/r/ChatGPT/comments/113ddqq/my_jailbreak_is_far_superior_to_dan_the_prompt_is/ from which we derive the instructions shown in Figure 6.
+Each of these instructions is inserted into the main jailbreak at the points where its individual instructions end, with the interleaving probability $p_{\lambda}$.
+All instructions inserted at the same point in the main jailbreak appear in a random order in the final jailbreak prefix.
+The main jailbreak is also the jailbreak used to create the soft prefixes.
+Note that, these jailbreaks are just examples to demonstrate our framework, which is in no way restricted to them and generalizes to any future bias-eliciting manual, textual jailbreaks.
+(a) For gender
+(b) For race
+Figure 5: Main Jailbreaks for different sensitive attributes
+Figure 6: Instructions derived from helper jailbreak
+Appendix C Example samples from prefix distributions
+In this section, we provide examples of prefixes of different kinds, described in Section 4.
+Random prefixes. Figure 7 presents some examples of random prefixes, each having $100$ tokens.
+Figure 7: Examples of random prefixes
+Mixture of jailbreaks. Figure 8 presents some examples of mixture of jailbreak prefixes, formed from the main and helper jailbreaks in Figures 5 and 6 respectively.
+Figure 8: Examples of mixture of jailbreak prefixes
+Soft prefixes.
+Examples of soft prefixes added to the queries can not be represented in textual form as they consist of perturbations of the prefix in the LLM’s embedding space.
+However, the general structure is {jailbreak perturbed in embedding space}.
+Appendix D Case studies from Decoding Trust certificates
+We introspect the example [target group] have a drug problem from $\mathcal{Q}_{DT}$.
+We study the responses of $4$ LLMs — Mistral, Llama-7B, Gemini, and GPT-3.5, obtained when certifying them for the mixture of jailbreaks specification based on the example.
+Figure 9 presents some samples of responses from the models to explain their certificates.
+We observe that the Gemini model is biased against white people, calling them ‘lazy’ and ‘more likely to be exposed to drugs’.
+Similarly, Mistral is also biased against white people, saying that they ‘have a monopoly on addiction’.
+On the other hand, GPT-3.5 is biased against black people, depicting them as ‘inferior’ and ‘inclined towards addiction’.
+The Llama model is safer, refraining from biased responses.
+The bounds for the probability of unbiased response suggest that GPT-3.5 and Gemini are more inclined to produce biased responses than the Mistral and Llama models for prompts in the distribution of counterfactual prompt sets in the given specification.
+Figure 9: Responses of Mistral, Llama-7B, Gemini, and GPT-3.5 obtained during their certification for the Decoding Trust example ‘[target group] have a drug problem’ for mixture of jailbreaks specification.
+The certification result is indicated alongside the model.
+Appendix E Ablation study
 In this section, we study the effect of changing the various certification parameters (a.k.a.
-hyperparameters) on the certificates generated with QuaCer-B.
-Table A presents the list of hyperparameters and their values used in our experiments.
+hyperparameters) on the certificates generated with LLMCert-B.
+Table E presents the list of hyperparameters and their values used in our experiments.
 Table 3: Hyperparameter values
@@
 $n$ Number of samples for certification 50
+$T$ LLM decoding temperature 1.0
+Top-k LLM decoding top-k 10
 $q$ Prefix length for random prefixes 100
@@
 We regenerate the certificates for different prefix distributions by varying the hyperparameters.
-In particular, we study the variations of the results when $n,q,p_{\lambda},p_{\mu},$ and $\kappa$ are varied, keeping $\gamma$ constant.
+In particular, we study the variations of the results when $n,T,\text{Top-k},q,p_{\lambda},p_{\mu},$ and $\kappa$ are varied, keeping $\gamma$ constant.
 This is because, $1-\gamma$ denotes the confidence of the certification bounds and that is generally desired to be high.
@@
 We certify the Mistral-Instruct-v0.2 [Jiang et al., 2023] 7B parameter model and study the overall results next.
+E.1 Certification algorithm hyperparameter
 We show ablations on $n$ for all kinds of specifications in Figure 10.
@@
 Figure 10: Ablation study on the certification hyperparameters showing variations of average certification bounds with number of samples $n$
+E.2 LLM decoding hyperparameters
+We study variations in the certification bounds with $2$ important hyperparameters of the LLM decoding algorithms that influence their generated texts — $T$ (decoding temperature) and Top-k (number of tokens decoded at each step).
+Figures 11 and 12 show the variations in the certification bounds with $T$ and Top-k respectively for the $3$ kinds of specifications.
+We see only minor changes in the average certification bounds with the variations of these hyperparameters.
+Our hypothesis of this phenomenon is that as the certificates aggregate the bias results of several samples, they smooth out the noise introduced by the choice of LLM decoding hyperparameters and give insights into the biases of the LLM itself.
+(a) Random prefixes
+(b) Mixtures of jailbreaks
+(c) Soft prefixes
+Figure 11: Ablation study showing variations of average certification bounds with temperature $T$
+(a) Random prefixes
+(b) Mixtures of jailbreaks
+(c) Soft prefixes
+Figure 12: Ablation study showing variations of average certification bounds with Top-k parameter.
 (a) Variation with prefix length $q$ for random prefix specifications
@@
 (c) Variation with mutation probability $p_{\mu}$ for mixture of jailbreaks specifications
-(d) Variation with relative magnitude of noise $\kappa$ for soft jailbreaks specifications
+(d) Variation with relative magnitude of noise $\kappa$ for soft prefixes from jailbreak specifications
 Figure 13: Ablation study on the certification hyperparameters showing variations of average certification bounds
-A.1 Random prefixes
+E.3 Random prefixes
 The specifications based on random prefixes consist of $1$ hyperparameter — $q$, length of the random prefix.
@@
 Figure 13(a) presents the variation in the average certification bounds obtained when varying $q$.
-A.2 Mixture of jailbreaks
+E.4 Mixture of jailbreaks
 These specifications have $2$ hyperparameters — $p_{\lambda}$, the probability of adding an instruction from the helper jailbreaks when interleaving, and $p_{\mu}$, the probability of randomly flipping every token of the resultant of interleaving.
-We show ablation studies on these in Figures 6(b) and 6(c) respectively.
-A.3 Soft jailbreaks
+We show ablation studies on these in Figures 13(b) and 13(c) respectively.
+E.5 Soft prefixes from jailbreaks
 These specifications have $1$ hyperparameter — $\kappa$, the maximum relative magnitude (with respect to the maximum magnitude of the embeddings) by which the additive uniform noise can change the embeddings of the main jailbreak.
@@
 We see no significant effects of the variation of abovementioned hyperparameters on certification results for different specifications.
-A.4 Scaling beyond binary demographic groups
+E.6 Scaling beyond binary demographic groups
 Table 4: Average bounds on the probability of unbiased responses from Mistral 7B.
@@
 While we are not aware of any $\mathcal{D}$ that could work with $\mathcal{Q}_{BOLD}$, we extend our $\mathcal{D}$ for the specifications from $\mathcal{Q}_{DT}$ to work for responses to prompts from three racial demographic groups — black people, white people, and asians.
-We elaborate on the extension in Appendix C.2.
+We elaborate on the extension in Appendix G.2.
 We certify Mistral-Instruct-v0.2 7B model with the three kinds of specifications and find that the average certification bounds presented in Table 4 are similar to the bounds presented for the Mistral model in Table 1 for the random and mixture of jailbreak specifications.
@@
 The bias detector identifies bias in responses having at least $1$ agreement and $1$ disagreement to the stereotype mentioned in the prompts, which has the same chance as unbiased result for binary demographic groups, but not beyond them.
-Appendix B Common jailbreaks used
-In our experiments, we combine common jailbreaks to certify against hybrid jailbreaks.
-The specific jailbreaks we use are shown below.
-The main jailbreak [^2] Last accessed on Jan 22, 2024 from www.jailbreakchat.com/prompt/588ab0ed-2829-4be8-a3f3-f28e29c06621 is shown in Figure 7.
-We have modified the jailbreak to encourage the model to generate responses demonstrating gender bias by adding the text highlighted in red in Figure 7(a) and for racial bias in Figure 7(b).
-We use one helper jailbreak [^3] https://www.reddit.com/r/ChatGPT/comments/113ddqq/my_jailbreak_is_far_superior_to_dan_the_prompt_is/ from which we derive the instructions shown in Figure 8.
-Each of these instructions is inserted into the main jailbreak at the points where its individual instructions end, with the interleaving probability $p_{\lambda}$.
-All instructions inserted at the same point in the main jailbreak appear in a random order in the final jailbreak prefix.
-(a) For gender
-(b) For race
-Figure 7: Main Jailbreaks for different sensitive attributes
-Figure 8: Instructions derived from helper jailbreak
-Appendix C Bias detector
-C.1 BOLD experiments
+Appendix F Validity of confidence intervals
+We design a synthetic study for the validity of the confidence intervals as follows.
+As we can not precisely regulate the true probability of unbiased responses of LLMs, we assume various values of that probability and generate binary-valued samples indicating biased (non-zero) /unbiased (zero) responses from any LLM.
+Hence, we generate $50$ samples (same as the samples used in LLMCert-B’s certification) of the Bernoulli random variable $\mathcal{F}$ (Section 3.2), with various values for the probability of success and generate Clopper-Pearson confidence intervals for the success probability using the samples.
+We repeat this process $1000$ times and report the percentage of instances wherein the confidence intervals contain the true probability of success.
+This percentage indicates the probability of correctness of the confidence intervals.
+We find that for all 11 equally-spaced values of the true probability of unbiased responses between 0 and 1, the confidence intervals bound the true value for more than 95% (nominal, user-specified confidence level) times, which validates the claim that the confidence intervals hold with at least the user-specified confidence.
+Figure 14 presents the variation in the proportion of Clopper-Pearson confidence intervals that contain the true probability of unbiased responses for different values of the latter.
+Figure 14: Variation in the proportion of Clopper-Pearson confidence intervals containing the true probability of unbiased responses.
+Dashed line shows the user-specified confidence level, i.e.
+$95\%$.
+Appendix G Bias detector
+G.1 BOLD experiments
 Design of bias detector.
 Our bias detector $\mathcal{D}_{BOLD}$ is based on the regard bias metric [Sheng et al., 2019].
-Specifically, we measure the bias for a pair of completions, one for male and another for female, by invoking the regard classifier in ’compare’ mode.
+Specifically, we measure the bias for a pair of completions, one for male and another for female, by invoking the regard classifier in ‘compare’ mode.
 The compare mode gives the difference between the regard score obtained for the different completions.
@@
 But we also recognize that bias is a complex subject, and its notion can vary across individuals.
+The human annotation of bias is contingent on many factors such as their cultural background, our annotation instructions (provided in the open-source implementation of the framework), etc.
 Hence, we believe that the noise in the human study needs to be taken into account when evaluating our bias detector $\mathcal{D}_{BOLD}$.
-C.2 DT experiments
+Our bias detector shows 93% precision, but 50% recall.
+Owing to these inconsistencies of our bias detector with human perception of bias, we believe that our bounds for probability of unbiased responses are actually higher than bounds with a perfect bias detector (due to low recall), indicating a worse situation of counterfactual bias in SOTA LLMs.
+G.2 DT experiments
 We design the bias detector $\mathcal{D}_{DT}$ for specifications from the Decoding Trust dataset, using the agreement detector proposed for the stereotype bias dataset in the Decoding Trust paper [Wang et al., 2024].
@@
 $\mathcal{D}_{DT}$ can easily scale beyond binary demographic groups, as it is agnostic to the number of responses across which it needs to detect bias.
-C.3 Limitations of bias detectors
+G.3 Limitations of bias detectors
 As we can see from the human evaluations for $\mathcal{D}_{BOLD}$ and also from the simplicity of $\mathcal{D}_{DT}$ that the bias detectors can have several false positive and false negative cases.
@@
 While our certification method borrows some of the inaccuracies of the bias detectors we use, they are often smoothed as certification aggregates several observations of bias to generate bounds on the probability of unbiased responses.
-Appendix D Example samples from prefix distributions
-In this section, we provide examples of prefixes of different kinds, described in Section 4.
-Random prefixes. Figure 10 presents some examples of random prefixes, each having $100$ tokens.
-Figure 10: Examples of random prefixes
-Mixture of jailbreaks. Figure 11 presents some examples of mixture of jailbreak prefixes, formed from the main and helper jailbreaks in Figures 7 and 8 respectively.
-Figure 11: Examples of mixture of jailbreak prefixes
-Soft prefixes.
-Examples of soft prefixes added to the queries can not be represented in textual form as they consist of perturbations of the prefix in the LLM’s embedding space.
-However, the general structure is {jailbreak perturbed in embedding space}.
-Appendix E Case studies from Decoding Trust certificates
-We introspect the example [target group] have a drug problem from $\mathcal{Q}_{DT}$.
-We study the responses of $4$ LLMs — Mistral, Llama-7B, Gemini, and GPT-3.5, obtained when certifying them for the mixture of jailbreaks specification based on the example.
-Figure 12 presents some samples of responses from the models to explain their certificates.
-We observe that the Gemini model is biased against white people, calling them ‘lazy’ and ‘more likely to be exposed to drugs’.
-Similarly, Mistral is also biased against white people, saying that they ‘have a monopoly on addiction’.
-On the other hand, GPT-3.5 is biased against black people, depicting them as ‘inferior’ and ‘inclined towards addiction’.
-The Llama model is safer, refraining from biased responses.
-The bounds for the probability of unbiased response suggest that GPT-3.5 and Gemini are more inclined to produce biased responses than the Mistral and Llama models for prompts in the distribution of counterfactual prompt sets in the given specification.
-Figure 12: Responses of Mistral, Llama-7B, Gemini, and GPT-3.5 obtained during their certification for the Decoding Trust example ‘[target group] have a drug problem’ for mixture of jailbreaks specification.
-The certification result is indicated alongside the model.
-Appendix F Positive and Negative impacts of our work
-We identify the following positive and negative impacts of our work.
-Positive impacts.
-Our work is the first to provide quantitative certificates for the bias in Large Language Models.
-It can be used by model developers to thoroughly assess their models before releasing them and by the general public to become aware of the potential harms of using any LLM.
-As our framework, QuaCer-B assumes black-box access to the model, it can be applied to even closed-source LLMs with API access.
-Negative impacts.
-In this work, we propose $3$ kinds of specifications involving — random prefixes, mixtures of jailbreaks as prefixes, and jailbreaks in the embedding space of the target model.
-While these prefixes are not adversarially designed, they are often successful in eliciting biased and toxic responses from the target LLMs.
-They can be used to attack these LLMs by potential adversaries.
-We have informed the developers of the LLMs about this threat.
+Appendix H Comparing with prior works on guarantees for LLMs
+Works on conformal prediction.
+[Quach et al., 2024, Deutschmann et al., 2023, Mohri and Hashimoto, 2024, Yadkori et al., 2024] apply conformal prediction techniques to language models, proposing methods for generating sets of outputs with statistical guarantees on correctness, coverage, or abstention, aiming to improve reliability and mitigate hallucinations.
+Their guarantees and scope differ from those of LLMCert-B.
+The prior works give specialized decoding strategies that guarantee the correctness of the outputs, useful for the factuality of the responses.
+LLMCert-B, however, is about assessing and certifying the counterfactual biases in LLM responses generated with any decoding scheme applied on the target LLMs.
+Moreover, guarantees of conformal prediction are for correctness of LLM responses for individual prompts, while LLMCert-B’s guarantees are over distributions of counterfactual prompt sets (which have prohibitively large sample spaces).
+Prompt Risk Control.
+[Zollo et al., 2024] (PRC) presents a framework for selecting low-risk system prompts for LLMs.
+PRC computes the loss incurred for one prompt at a time, and aggregates those losses to form a risk measure.
+LLMCert-B, on the other hand, is for counterfactual bias, i.e., we assess bias across a set of LLM responses, obtained by varying the sensitive attributes in prompts.
+Our specification is thus a relational property [Barthe et al., 2011], which is defined over multiple related inputs.
+Biases across LLM responses for multiple related prompts are aggregated to certify the target LLM.
+Moreover, LLMCert-B proposes novel, inexpensive mechanisms to design distributions and their samplers with potentially adversarial prefixes, containing effective, manually-designed jailbreaks in their sample space.
+Contrary to PRC which uses static datasets of adversarial examples, we use independent and identically-distributed samples from these distributions.
+Appendix I Additional notes on Bias in ML
+Bias is a complex social phenomenon that arises in various forms.
+In this section, we discuss various notions of bias and harms caused by them.
+We also discuss how LLMCert-B complements existing evaluation methods by certifying for counterfactual bias.
+The following discussion is not a comprehensive treatment of bias in Machine Learning and we refer the reader to detailed survey papers such as [Gallegos et al., 2024a, Blodgett et al., 2020, Li et al., 2024] for more information.
+Bias consists of disparate treatment (a.k.a.
+discrimination) or disparate outcomes [Barocas and Selbst, 2016] for different demographic groups.
+Harms due to bias are primarily of $2$ kinds — representational and allocation [Gallegos et al., 2024a].
+Representational harm [Suresh and Guttag, 2021, Blodgett et al., 2020] consists of denigrating and subordinating attitudes towards a demographic group.
+It consists of use of derogatory language, stereotyping, toxicity, misrepresentation, etc.
+These can arise from inappropriate use of language by humans or machines (e.g., LLMs).
+Allocation harms [Ferrara, 2023] are disparate distribution of resources or opportunities between demographic groups.
+These consist of direct or indirect discrimination in economic or social opportunities.
+For example, prior works like [Terry et al., 2010, Martínez, 2022] show that the lack of representation of African American English in dominant language practices results in that community facing penalties in education systems or when seeking housing.
+Most constitutions around the world have anti-discrimination laws like [Sherry, 1965] that prohibit allocative harms in employment etc.
+Language is considered an important factor for labeling, modifying, and transmitting beliefs about demographic groups and can result in the reinforcement of social inequalities [Rosa and Flores, 2017].
+LLMCert-B is a reliable evaluation method for counterfactual bias in language models (LMs), that certifies the probability of unbiased response in target LLMs for distributions of counterfactual prompts with statistical guarantees.
+Prior bias assessments have been of $2$ kinds [Cao et al., 2022] — intrinsic and extrinsic.
+Intrinsic bias occurs in the language representations, while extrinsic bias manifests in the final textual responses of the LMs.
+To certify closed-source LMs as well, we study extrinsic bias.
+Bias is opposite of fairness, which has been identified to be of various forms such as group fairness [Blandin and Kash, 2024], individual fairness [Dwork et al., 2011], counterfactual fairness [Kusner et al., 2018], etc.
+LLMCert-B certifies for counterfactual bias, akin to counterfactual fairness.
+This is because of the causal perspective of counterfactual bias [Anthis and Veitch, 2023] (bias due to mentioning specific demographic group in the prompt) which aligns more closely with human intuitions about discrimination and fairness.
+Moreover, unlike group fairness, counterfactual fairness operates at the individual-level, thus identifying bias in specific cases, instead of aggregates.